### 15 Mar 2024 v3

# Low level definability above large cardinals

Abstract
We study some connections between definability in generalized descriptive set theory and large cardinals, particularly measurable cardinals and limits thereof, working in ZFC. We show that if $\kappa$ is a limit of measurable cardinals then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ wellorder of a subset of $\mathcal{P}(\kappa)$ of length $\geq\kappa^{+}$; this answers a question of Lücke and Müller. However, in $M_{1}$, the minimal proper class mouse with a Woodin cardinal, for every uncountable cardinal $\kappa$ which is not a limit of measurables, there is a $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ good wellorder of $\mathcal{H}_{\kappa^{+}}$. If $\kappa$ is a limit of measurables then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$, and if also $\operatorname{cof}(\kappa)>\omega$ then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ almost disjoint family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$. However, relative to the consistency of large cardinals, $\Pi_{1}(\{\kappa\})$ mad families and maximal independent families $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ can exist, when $\kappa$ is a limit of measurables, and even more. We also examine some of the features of $L[U]$, and answer another question of Lücke and Müller, showing that if $\kappa$ is a weakly compact cardinal such that every $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ subset of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$ contains the range of a perfect function, then there is an inner model satisfying “there is a weakly compact limit of measurable cardinals”.
1 Introduction
In this paper we study the definability of particular kinds of subsets of generalized Baire space ${{}^{\kappa}}\kappa$, for uncountable cardinals $\kappa$, in the presence of large cardinals. The kinds of sets considered are:
  1. wellorders of subsets of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$, and in particular, good wellorders of $\mathcal{P}(\kappa)$,
  2. almost disjoint families at $\kappa$ of cardinality $>\kappa$,
  3. subsets of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$, which have no perfect subset,
  4. and some brief remarks on ultrafilters over $\kappa$ and on the club filter at $\kappa$.
The work relates particularly to that in Lücke and Müller [9], Lücke, Schindler and Schlicht [10], and Lücke and Schlicht [11]; in particular, we answer some of the questions posed in [9].
Given infinite cardinals $\mu\leq\nu$, say $\nu$ is $\mu$-closed iff $\forall\alpha<\nu\ [\alpha^{\mu}<\nu]$. Given infinite cardinals $\mu\leq\kappa$, say $\kappa$ is $\mu$-steady iff there is a cardinal $\nu\leq\kappa$ such that $\nu$ is $\mu$-closed, $\operatorname{cof}(\nu)\neq\mu$ and $\kappa\in\{\nu,\nu^{+}\}$. So if $\kappa$ is $\mu$-steady then $\kappa>\mu^{+}$. Note that if $\nu$ is $\mu$-closed and $\operatorname{cof}(\nu)<\mu$ then $\nu^{+}$ is $\mu$-steady but non-$\mu$-closed, as $\nu^{\mu}\geq\nu^{+}$.
Regarding wellorders, first, in §3.1, we consider cardinals $\kappa$ such that there is a measurable cardinal $\mu<\kappa$ and $\kappa$ is $\mu$-steady. Under these assumptions, in Theorem 3.2, we establish some restrictions on wellordered subsets of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$ which are $\Sigma_{1}$ definable in certain parameters; these results are minor refinements of results in [9] and [11]. The main new fact here is that there is no $\Sigma_{1}(V_{\mu}\cup\{\kappa\})$ injection from $\kappa^{+}$ into $\mathcal{P}(\kappa)$, even if $\operatorname{cof}(\kappa)=\omega$; this was proved under the added assumption that $\operatorname{cof}(\kappa)>\omega$ in [9, Theorem 7.1]. In fact, we show that there is no $\Sigma_{1}(V_{\mu}\cup\{\kappa\})$ set $f\subseteq\kappa^{+}\times\mathcal{P}(\kappa)$ such that for some club $C\subseteq\kappa^{+}$, $f\!\upharpoonright\!C:C\to\mathcal{P}(\kappa)$ is an injective function.
In §3.2 we consider cardinals $\kappa$ which are limits of measurables. Here we answer [9, Questions 10.3, 10.4], by showing that in this context, there is no set $D\subseteq\mathcal{P}(\kappa)$ and wellorder $<^{*}$ of $D$ such that $D$ has cardinality $>\kappa$ and $D,<^{*}$ are $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$-definable. (Lücke and Müller already established some results in this direction in [9, Theorem 1.4, Corollary 7.4].) We also strengthen their results [9, Theorem 1.1, Corollary 7.4(ii)], as explained in detail in §3.2.
In §3.3, we prove that if $j:V\to M$ is an $I_{2}$-embedding and $\kappa$ is the limit its critical sequence, then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\mathcal{H}_{\kappa}\}\cup\mathrm{OR})$ wellorder of a subset of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$.
In §4, we consider the situation in the model $L[U]$ for one measurable cardinal $\mu$. We show that in $L[U]$, for every uncountable cardinal $\kappa$:
  1. (i)
    there is a stationary-co-stationary set $d\subseteq\kappa^{+}$ and an injective function $f:d\to\mathcal{P}(\kappa)$ such that $d,f$ are both $\Sigma_{1}(\{\kappa\})$, and
  2. (ii)
    there is a $\Sigma_{1}(\{\kappa\})$ set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$ such that there is no perfect embedding $\iota:{{}^{\operatorname{cof}(\kappa)}}\kappa\to{{}^{\kappa}}\kappa$ with $\mathrm{rg}(\iota)\subseteq D$.
(In case $\kappa\leq\mu^{+}$ or $\kappa$ is non-$\mu$-steady, these things already follow easily from the results in [10]. Part (ii) relates to [9, Theorem 1.2], by which it was already known hat if $\kappa$ is singular then there must be some such $D$ which is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$. But that result did not deal with the case that $\kappa$ is regular.)
In Theorem 5.1, we establish the existence, in $M_{1}$ (the minimal proper class mouse with a Woodin cardinal), of a $\Sigma_{1}(\mathcal{H}^{M_{1}}_{\kappa}\cup\{\kappa\})$ good wellorder of $\mathcal{H}^{M_{1}}_{\kappa^{+M_{1}}}$, for all uncountable cardinals $\kappa$ which are not limits of measurables in $M_{1}$.
In §6 we answer [9, Question 10.1], which asked whether, if $\kappa$ is weakly compact and $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ has a certain perfect embedding property, there must be a (proper class) inner model satisfying “there is a weakly compact limit of measurable cardinals”. In Theorem 6.1 we show the answer is “yes” (the “perfect embedding property” is clarified in the theorem’s statement).
In §7, we consider almost disjoint families. Let $\kappa$ be an infinite cardinal. Recall that an almost disjoint family at $\kappa$ is a set $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ such that for all $A\in\mathscr{F}$, $A$ is unbounded in $\kappa$, and for all $A,B\in\mathscr{F}$ with $A\neq B$, $A\cap B$ is bounded in $\kappa$. We usually drop the phrase “at $\kappa$”, as $\kappa$ will be clear from context. An almost disjoint family $\mathscr{F}$ is called maximal if for all unbounded $C\subseteq\kappa$, there is $A\in\mathscr{F}$ such that $A\cap C$ is unbounded in $\kappa$. A mad family just means a maximal almost disjoint family. It was shown by Adrian Mathias in [12, Corollary 4.7] that there is no $\utilde{\Sigma}^{1}_{1}$ infinite mad family at $\omega$, and shown by Arnold Miller in [13, Theorem 8.23] that if $V=L$ then there is a $\Pi^{1}_{1}$ infinite mad family at $\omega$. In this section we establish variants of these results, considering almost disjoint families at cardinals $\kappa>\omega$. Lücke and Müller already proved an analogue [9, Theorem 1.3] of Mathias’ result, showing that if $\kappa$ is an iterable cardinal which is a limit of measurables then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ almost disjoint family $\mathscr{F}$ at $\kappa$ such that $\mathscr{F}$ has cardinality $>\kappa$, and hence no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ mad family $\mathscr{F}$ at $\kappa$ such that $\mathscr{F}$ has cardinality $\geq\kappa$. We extend this, showing in Theorem 7.1 that if $\kappa$ a limit of measurable cardinals then:
  1. there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\geq\kappa$, and
  2. if $\operatorname{cof}(\kappa)>\omega$ then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ almost disjoint family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$.
In the other direction, we establish an analogue of Miller’s result, and in so doing, partially address [9, Question 10.5(ii)], which asked whether sufficiently strong large cardinal properties of $\kappa$ imply that there is no $\Pi_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ almost disjoint family of cardinality $>\kappa$. We will show in Theorem 7.2 that it is in fact consistent relative to large cardinals that $\kappa$ is a regular cardinal, with large cardinal properties up to a Woodin limit of Woodin cardinals, and there is a $\Pi_{1}(\{\kappa\})$ mad family of cardinality $>\kappa$.
In §8 we establish in Theorem 8.2 an analogue of Theorem 7.2, giving the consistency relative to large cardinals of cardinals $\kappa$ with large cardinal properties, together with the existence of $\Pi_{1}(\{\kappa\})$ maximal independent families $\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$.
In §9 we make a simple observation on the definability of ultrafilters over $\kappa$ which are limits of measurables, and one on the definability of the club filter on regular $\kappa$ such that $\kappa>\mu$ for some measurable $\mu$.
Finally in §10 we consider a somewhat different theme. We adapt a result of Schlicht [15, Theorem 2.19], proving the relative consistency of the theory ZFC + “$\kappa$ is $\kappa^{+}$-supercompact” + “for every $X\subseteq\mathcal{P}(\kappa)$ with $X\in\mathrm{HOD}_{V_{\kappa+1}}$ and $X$ of cardinality $>\kappa$, we have that $X$ has a perfect subset and $X$ is not a wellorder. (The methods involved in this adaptation are some standard forcing techniques.)
In §11 we collect some questions together.
1.1 Notation
ZFC is the background theory for the paper.
For classes $X,Y$, we say $Y$ is $\Sigma_{n}(X)$ (or $Y$ is $\Sigma_{n}(X)$-definable) iff there is a $\Sigma_{n}$ formula $\varphi$ and elements $p_{0},\ldots,p_{n-1}\in X$ such that for all $y$, we have $y\in Y$ iff $\varphi(p_{0},\ldots,p_{n-1},y)$. Likewise for $\Pi_{n}(X)$. And $Y$ is $\Delta_{n}(X)$ iff it is both $\Sigma_{n}(X)$ and $\Pi_{n}(X)$.
For an ordinal $\eta$ let $[\eta]^{\omega}_{\uparrow}$ denote the set of all subsets of $\eta$ of ordertype $\omega$.
Recall that $\mu$-closed and $\mu$-steady were defined at the beginning of §1.
By a measure, we mean a countably complete non-principal ultrafilter, and by a normal measure, we mean a normal $\kappa$-complete non-principal ultrafilter on an uncountable cardinal $\kappa$.
For an uncountable cardinal $\kappa$, $\mathcal{H}_{\kappa}$ denotes the set of all sets hereditarily of cardinality $<\kappa$.
The universe of a structure $M$ is denoted $\left\lfloor M\right\rfloor$.
Recall from [25] that a premouse $M$ has form $M=(\mathcal{J}_{\alpha}[\mathbb{E}],\mathbb{E},F)$ where $\alpha$ is a limit ordinal or $\alpha=\mathrm{OR}$, $\mathbb{E}$ is a sequence of (partial) extenders with certain properties, and either $F=\emptyset$ or $F$ is an extender over $\mathcal{J}_{\alpha}[\mathbb{E}]$ with certain properties. We write $\mathbb{E}^{M}=\mathbb{E}$, $F^{M}=F$ and $\mathbb{E}_{+}^{M}=\mathbb{E}^{M}\ \widehat{\ }\ \left<F^{M}\right>$. We say $F$ is the active extender of $M$, and say $M$ is active if $F\neq\emptyset$. For limit ordinals $\beta\leq\mathrm{OR}^{M}$, we write $M|\beta$ for the initial segment of $M$ of ordinal height $\beta$, including the extender indexed at $\beta$, if there is one; that is, $M|\beta=(\mathcal{J}_{\beta}[\mathbb{E}^{M}],\mathbb{E}^{M}\!\upharpoonright\!\beta,E)$ where $E=\mathbb{E}_{+}^{M}(\beta)$. We write $M||\beta=(M|\beta)^{\mathrm{pv}}$ for $(\mathcal{J}_{\beta}[\mathbb{E}^{M}],\mathbb{E}^{M}\!\upharpoonright\!\beta,\emptyset)$. If $\beta\in\mathrm{OR}^{M}$ is an $M$-cardinal then $\mathbb{E}^{M}(\beta)=\emptyset$. We write $<^{M}$ for the standard order of constructibility of $M$ (this is a wellorder of $\left\lfloor M\right\rfloor$, and ${<^{M||\beta}}={<^{M|\beta}}$ is an initial segment of $<^{M}$, for all $\beta\leq\mathrm{OR}^{M}$). Actually we officially use the definition of premouse of [19, §1.1], and in particular, allow extenders of superstrong type to appear in $\mathbb{E}_{+}^{M}$ (though these will only be relevant for some of the work).
We also consider $L[U]$ (where $U$ is an filter over some $\kappa$ such that $L[U]\models$$U$ is a normal measure”) to be a premouse in the above sense, so $\mathbb{E}^{L[U]}$ includes $U$ as its ultimate element, but also includes many partial measures prior to this (both on $\kappa$ and unboundedly many ordinals $<\kappa$).
2 Linear iterations of measures
In this section we collect some facts regarding linear iterations of measures which we will need. The material in §2.1 is standard, and is very similar to some material in [21].
2.1 Iterations of a single measure
2.1 Definition.
For a measure $U$ on a measurable cardinal $\mu$ let $\mathcal{T}_{U}=\left<M_{\alpha},U_{\alpha}\right>_{\alpha\in\mathrm{OR}}$ be the length $\mathrm{OR}$ iteration of $V$ via $U$ and its images. That is, $M_{0}=V$, $U_{0}=U$, $M_{\alpha+1}=\mathrm{Ult}(M_{\alpha},U_{\alpha})$, and letting $i_{\alpha\beta}:M_{\alpha}\to M_{\beta}$ be the iteration map, $U_{\alpha+1}=i_{0,\alpha+1}(U)=i_{\alpha,\alpha+1}(U_{\alpha})$, and for limit $\lambda$, $M_{\lambda}$ is the direct limit of the earlier $M_{\alpha}$ under these iteration maps. We also write $M^{\mathcal{T}_{U}}_{\alpha}=M_{\alpha}$, $i^{\mathcal{T}_{U}}_{\alpha\beta}=i_{\alpha\beta}$, etc.
In the following lemma, recall that a cardinal $\kappa$ is $\mu$-closed iff $\alpha^{\mu}<\kappa$ for all $\alpha<\kappa$, and $\mu$-steady iff there is a cardinal $\nu\leq\kappa$ such that $\nu$ is $\mu$-closed, $\operatorname{cof}(\nu)\neq\mu$, and $\kappa\in\{\nu,\nu^{+}\}$.
2.2 Lemma.
Assume ZFC and let $\mu<\kappa$ be cardinals such that $\mu$ is measurable and $\kappa$ is $\mu$-closed or $\mu$-steady. Let $U$ be a $\mu$-complete measure on $\mu$. Then:
  1. 1.
    $i^{\mathcal{T}_{U}}_{0\lambda}``\kappa\subseteq\kappa$ for each $\lambda<\kappa$.
  2. 2.
    There are unboundedly many $\mu^{\prime}<\kappa$ such that $i^{\mathcal{T}_{U}}_{0\mu^{\prime}}(\mu)=\mu^{\prime}$ and $i^{\mathcal{T}_{U}}_{0\lambda}(\mu^{\prime})=\mu^{\prime}$ for all $\lambda<\mu^{\prime}$.
Proof.
Part 1: Let $\alpha<\kappa$. If $\kappa$ is $\mu$-closed then $\mathrm{card}(i^{\mathcal{T}_{U}}_{0\lambda}(\alpha))\leq\mathrm{card}(\alpha)^{\mu}\cdot(\mathrm{card}(\lambda+1))<\kappa$. Suppose $\kappa$ is non-$\mu$-closed. Then $\kappa$ is $\mu$-steady, and so $\kappa=\nu^{+}$ where $\nu$ is $\mu$-closed and $\operatorname{cof}(\nu)\neq\mu$ but $\nu^{\mu}\geq\kappa$. Note then that $\operatorname{cof}(\nu)<\mu$, but then $\mathrm{card}(i^{\mathcal{T}_{U}}_{0\lambda}(\nu)\leq\nu\cdot\lambda=\nu$, since to form $i^{\mathcal{T}_{U}}_{0\lambda}(\nu)$, it suffices to consider only functions $f:[\mu]^{<\omega}\to\nu$ which are bounded in $\nu$, of which there are only $\nu$-many. So $i^{\mathcal{T}_{U}}_{0\lambda}(\nu)<\nu^{+}=\kappa$. But then $i^{\mathcal{T}_{U}}_{0\lambda}(\alpha)<\kappa$ also.
Part 2: Fix $\lambda<\kappa$; we want to find $\mu^{\prime}\in[\lambda,\kappa)$ with the right properties. Let $\mu_{0}=i^{\mathcal{T}_{U}}_{0,\lambda+1}(\mu)$ (so $\lambda+1<\mu_{0}$), and given $\mu_{n}$, let $\mu_{n+1}=i^{\mathcal{T}_{U}}_{0\mu_{n}}(\mu)$. Then $\mu_{n}<\mu_{n+1}<\kappa$. Let $\mu^{\prime}=\sup_{n<\omega}\mu_{n}$. Some straightforward cardinal arithmetic shows that $\mu^{\prime}<\kappa$ (even if $\operatorname{cof}(\kappa)=\omega$), and note that $\mu^{\prime}$ works. ∎
2.3 Definition.
For an ordinal $\alpha$, the eventual ordertype $\mathrm{eot}(\alpha)$ of $\alpha$ is the least ordinal $\eta$ such that for some $\beta<\alpha$, we have $\alpha=\beta+\eta$.
2.4 Definition.
Let $U$ be a measure. Given a set $X$ and ordinals $\alpha\leq\beta$ with $X\in M^{\mathcal{T}_{U}}_{\gamma}$ for all $\gamma<\beta$, we say that $X$ is $U$-$[\alpha,\beta)$-stable iff for all $\gamma\in[\alpha,\beta)$ we have $i^{\mathcal{T}_{U}}_{\alpha\gamma}(X)=X$. If $U$ is determined by context, we just say $[\alpha,\beta)$-stable, and often we will have a cardinal $\kappa$ also fixed, and then ($U$)-$\alpha$-stable means ($U$)-$[\alpha,\kappa)$-stable.
A well-known fact is the following:
2.5 Fact.
Let $U$ be a measure. Let $\eta$ be a limit ordinal and $\xi$ any ordinal. Then there is $\alpha<\eta$ such that $\xi$ is $U$-$[\alpha,\eta)$-stable. (Otherwise observe that $M^{\mathcal{T}}_{\eta}$ is illfounded.)
2.6 Definition.
Let $U$ be a measure and $\eta\in\mathrm{OR}$ be a limit ordinal. Then *$*_{\eta,U}:\mathrm{OR}\to\mathrm{OR}$ denotes the map $\alpha\mapsto\alpha^{*}$ where for ordinals $\alpha$ we define $\alpha^{*}=i^{\mathcal{T}_{U}}_{\beta\eta}(\alpha)$ for any/all $\beta<\eta$ such that $\alpha$ is $[\beta,\eta)$-stable. We just write *$*_{\eta}$ if $U$ is clear from context, and just write $\alpha^{*}$ for *$*_{\eta}(\alpha)$ if $\eta$ is also clear.
A straightforward calculation shows the following fact:
2.7 Lemma.
Let $U,\eta$ be as in Definition 2.6. Let $\tau=\mathrm{eot}(\eta)$ and $\tau^{\prime}=*_{\eta}(\tau)$. Then for all $\alpha\in\mathrm{OR}$, we have
*$*_{\eta}(\alpha)=i^{\mathcal{T}_{U}}_{\eta,\eta+\tau^{\prime}}(\alpha)=i^{\mathcal{T}_{U_{\eta}}}_{0\tau^{\prime}}(\alpha),$
and hence, *$*_{\eta}$ is definable over $M^{\mathcal{T}_{U}}_{\eta}$ from the parameters $U_{\eta}$ and $\tau^{\prime}$.
2.8 Lemma.
Let $U$ be a measure. Let $\eta$ be a limit ordinal and $\tau=\mathrm{eot}(\eta)$. Let $A\subseteq\mathrm{OR}$ be a set. Then $A\in M^{\mathcal{T}_{U}}_{\eta}$ iff $A\in\bigcap_{\alpha<\eta}M^{\mathcal{T}_{U}}_{\alpha}$ and there is $\beta<\eta$ such that $A$ is $U$-$[\beta,\eta)$-stable.
Proof.
Suppose $A\in M^{\mathcal{T}_{U}}_{\eta}$. Let $\beta<\eta$ be such that $A\in\mathrm{rg}(i^{\mathcal{T}_{U}}_{\beta\eta})$, $\beta+\tau=\eta$ and $\tau$ is $[\beta,\eta)$-stable. Let us show that $A$ is $[\beta,\eta)$-stable. For $\gamma\in[\beta,\eta)$ let $A=i^{\mathcal{T}_{U}}_{\gamma\eta}(A_{\gamma})$. Then note that for such $\gamma$, we have
$M^{\mathcal{T}_{U}}_{\gamma}\models\text{``}A=i^{\mathcal{T}_{U_{\gamma}}}_{0\tau}(A_{\gamma})\text{''}.$ ” . (1)
But then applying $i^{\mathcal{T}_{U}}_{\gamma\delta}$ to this when $\delta\in[\gamma,\eta)$, and letting $A^{\prime}=i^{\mathcal{T}_{U}}_{\gamma\delta}(A)$, we have
$M^{\mathcal{T}_{U}}_{\delta}\models\text{``}A^{\prime}=i^{\mathcal{T}_{U_{\delta}}}_{0\tau}(A_{\delta})\text{''.}$ ”. (2)
But then comparing lines (1) (where we can now replace $\gamma$ with $\delta$) and (2), we have $A^{\prime}=A$.
Now suppose that $A\in\bigcap_{\alpha<\eta}M^{\mathcal{T}_{U}}_{\alpha}$ and there is $\beta<\eta$ such that $A$ is $[\beta,\eta)$-stable. We show that $A\in M^{\mathcal{T}_{U}}_{\eta}$. We may take $\beta<\eta$ such that $A$ is $[\beta,\eta)$-stable, $\eta=\beta+\tau$ and $\tau$ is $[\beta,\eta)$-stable. Let $A^{*}=i^{\mathcal{T}_{U}}_{\beta\eta}(A)$, so $A^{*}=i^{\mathcal{T}_{U}}_{\gamma\eta}(A)$ for all $\gamma\in[\beta,\eta)$. Then for all $\alpha\in\mathrm{OR}$, note that $\alpha\in A$ iff $\alpha^{*}\in A^{*}$. Since $A^{*}\in M^{\mathcal{T}_{U}}_{\eta}$, by Lemma 2.7, we can deduce that $A\in M^{\mathcal{T}_{U}}_{\eta}$. ∎
2.2 Iterations of multiple measures
2.9 Definition.
Say that a linear iteration $\mathcal{T}$ on $V$ [^2] Note that we are no longer restricting to the kinds of iterations considered in 2.1, in that it need not be that $E^{\mathcal{T}}_{\alpha}=i^{\mathcal{T}}_{0\alpha}(U)$ for some fixed $U$, nor that $E^{\mathcal{T}}_{\alpha}\in\mathrm{rg}(i^{\mathcal{T}}_{0\alpha})$, nor even that $\mathrm{cr}(E^{\mathcal{T}}_{\alpha})\in\mathrm{rg}(i^{\mathcal{T}}_{0\alpha})$. is a linear iteration of measures iff $M^{\mathcal{T}}_{\alpha}\models$$E^{\mathcal{T}}_{\alpha}$ is a measure (not necessarily normal)” for all $\alpha+1<\mathrm{lh}(\mathcal{T})$. For a linear iteration of measures, say that $\mathcal{T}$ is above $\mu$ if $\mu\leq\mathrm{cr}(i^{\mathcal{T}})$ (equivalently, $M^{\mathcal{T}}_{\alpha}\models$$E^{\mathcal{T}}_{\alpha}$ is $\mu$-complete” for all $\alpha+1<\mathrm{lh}(\mathcal{T})$), and based on $V_{\delta}$ iff $E^{\mathcal{T}}_{\alpha}\in V_{i^{\mathcal{T}}_{0\alpha}(\delta)}^{M^{\mathcal{T}}_{\alpha}}$ for all $\alpha+1<\mathrm{lh}(\mathcal{T})$.
The following lemma is just a very slight variant on a well-known fact (see Kunen [8], and very related calculations in [24, Lemma 4.5] and [17, Lemma 3.17***], for example). We give details for self-containment.
2.10 Lemma.
Assume ZFC. Let $\mathcal{T}$ and $\mathcal{U}$ be a linear iterations of measures on $V$, of successor lengths $\alpha+1$ and $\beta+1$ respectively. Suppose there is an inaccessible cardinal $\mu$ such that $\alpha<\mu$, $\mathcal{T}$ is based on $V_{\mu}$, and $\mathcal{U}$ is above $\mu$. Let $j=i^{\mathcal{U}}:V\to M=M^{\mathcal{U}}_{\infty}$ be the iteration map. Let $V^{\prime}=M^{\mathcal{T}}_{\infty}$ and $\mathcal{U}^{\prime}=i^{\mathcal{T}}\mathcal{U}$ (the copy of $\mathcal{U}$ under $i^{\mathcal{T}}$). [^3] See [25, §4.1], though that version is much more general. Define $\pi_{0}=i^{\mathcal{T}}$. In general we set $E^{\mathcal{U}^{\prime}}_{\alpha}=\pi_{\alpha}(E^{\mathcal{U}}_{\alpha})$, then define $\pi_{\alpha+1}:M^{\mathcal{U}}_{\alpha+1}\to M^{\mathcal{U}^{\prime}}_{\alpha+1}$ as the natural map with $\pi_{\alpha+1}\circ i^{\mathcal{U}}_{\alpha,\alpha+1}=i^{\mathcal{U}^{\prime}}_{\alpha,\alpha+1}\circ\pi_{\alpha}$ (see [25, Lemma 4.2]), and for limit $\lambda$, $\pi_{\lambda}:M^{\mathcal{U}}_{\lambda}\to M^{\mathcal{U}^{\prime}}_{\lambda}$ similarly commutes with the earlier $\pi_{\alpha}$’s and iteration maps. We set $\mathrm{lh}(\mathcal{U}^{\prime})=\mathrm{lh}(\mathcal{U})$. If $\mathrm{lh}(\mathcal{U})<\mathrm{cr}(i^{\mathcal{T}})$ then $\mathcal{U}^{\prime}=i^{\mathcal{T}}(\mathcal{U})$ and $\pi_{\alpha}=i^{\mathcal{T}}\!\upharpoonright\!M^{\mathcal{U}}_{\alpha}$, but if $\mathrm{lh}(\mathcal{U})\geq\mathrm{cr}(i^{\mathcal{T}})$ then this breaks down. Let $M^{\prime}=M^{\mathcal{U}^{\prime}}_{\infty}$ and $j^{\prime}=i^{\mathcal{U}^{\prime}}:V^{\prime}\to M^{\prime}$ be the iteration map. Then $j(V^{\prime})=M^{\prime}$ and $j\!\upharpoonright\!V^{\prime}=j^{\prime}$.
Proof.
Let us first consider the case that $\mathcal{U}$ has finite length, so is just a finite iteration of measures, each of which are $\mu$-complete. Thus, there is a $\mu$-complete measure $U$ on some $\nu\geq\mu$ such that $M^{\mathcal{U}}_{\infty}=\mathrm{Ult}(V,U)$ and $i^{\mathcal{U}}=i^{V}_{U}$. Moreover, since $\mathcal{U}$ is finite, $\mathcal{U}^{\prime}=i^{\mathcal{T}}(\mathcal{U})$; let $U^{\prime}=i^{\mathcal{T}}(U)$, a measure on $\nu^{\prime}=i^{\mathcal{T}}(\nu)$, so $M^{\mathcal{U}^{\prime}}_{\infty}=\mathrm{Ult}(V^{\prime},U^{\prime})$ and $i^{\mathcal{U}^{\prime}}=i^{V^{\prime}}_{U^{\prime}}$.
Now we define a map
$\pi:\mathrm{Ult}(V^{\prime},U^{\prime})\to j(V^{\prime})$
as follows. For $f\in V^{\prime}$ with $f:\nu^{\prime}\to V^{\prime}$, let
$\pi([f]^{V^{\prime}}_{U^{\prime}})=[f\circ k]^{V}_{U}$
where $k=i^{\mathcal{T}}\!\upharpoonright\!\nu:\nu\to i^{\mathcal{T}}(\nu)$.
Let us verify that this is well-defined. Let $f,g\in V^{\prime}$ with $f,g:\nu^{\prime}\to V^{\prime}$ and suppose that $[f]^{V^{\prime}}_{U^{\prime}}=[g]^{V^{\prime}}_{U^{\prime}}$, so
$X=_{\mathrm{def}}\{\alpha<\nu\bigm{|}f(\alpha)=g(\alpha)\}\in U^{\prime}.$
Note that there is $Y\in U$ such that $i^{\mathcal{T}}(Y)\subseteq X$. But then for each $\alpha\in Y$, we have $i^{\mathcal{T}}(\alpha)\in X$, and so $(f\circ k)(\alpha)=f(i^{\mathcal{T}}(\alpha))=g(i^{\mathcal{T}}(\alpha))=(g\circ k)(\alpha)$, which shows that $[f\circ k]^{V}_{U}=[g\circ k]^{V}_{U}$, as desired.
So $\pi$ is well-defined, and similarly, $\pi:\mathrm{Ult}(V^{\prime},U^{\prime})\to j(V^{\prime})$ is elementary.
We now claim that $\pi$ is surjective. For let $h:\nu\to V^{\prime}$. Note that since $\mu$ is inaccessible, we can fix $\delta<\mu$ such that $\mathcal{T}$ only uses measures from $V_{\delta}$ and its images. For each $\alpha<\nu$, let $(f_{\alpha},b_{\alpha})$ be such that $f_{\alpha}:\delta^{<\omega}\to V$ and $b_{\alpha}\in i^{\mathcal{T}}(\delta)^{<\omega}$ and $h(\alpha)=i^{\mathcal{T}}(f_{\alpha})(b_{\alpha})$. By $\mu$-completeness, and since $i^{\mathcal{T}}(\delta)<\mu$, we can fix $Y\in U$ and $b$ such that for all $\alpha\in Y$, we have $b_{\alpha}=b$. Now define $\widetilde{h}:\delta^{<\omega}\to V$ by setting $\widetilde{h}(x)$ to be the function such that $\widetilde{h}(x):\nu\to V$ and $\widetilde{h}(x)(\alpha)=f_{\alpha}(x)$. Let $f=i^{\mathcal{T}}(\widetilde{h})(b)$. Then note that $f\in V^{\prime}$ and $f:\nu^{\prime}\to V^{\prime}$ and for all $\alpha\in Y$, we have
$(f\circ k)(\alpha)=f(i^{\mathcal{T}}(\alpha))=i^{\mathcal{T}}(f_{\alpha})(b)=h(\alpha).$
Therefore $\pi([f]^{V^{\prime}}_{U^{\prime}})=[h]^{V}_{U}$, so $\pi$ is surjective, as desired.
This shows that $\pi$ is in fact an isomorphism. So we have that $j(V^{\prime})=\mathrm{Ult}(V^{\prime},U^{\prime})$. Finally, the fact that $j\!\upharpoonright\!V^{\prime}=j^{\prime}$ is just because for $x\in V^{\prime}$, we have $j^{\prime}(x)=[c_{x}]^{V^{\prime}}_{U^{\prime}}$ where $c_{x}:\nu^{\prime}\to V^{\prime}$ is the constant function $c_{x}(\alpha)=x$, and so
$\pi(j^{\prime}(x))=\pi([c_{x}]^{V^{\prime}}_{U^{\prime}})=[c_{x}\circ k]^{V}_{U}=j(x),$
and therefore $j^{\prime}(x)=j(x)$, as desired.
To generalize to the case that $\mathcal{U}$ has infinite length, just use the fact that $\mathcal{U}$ is the direct limit of iterations $\bar{\mathcal{U}}$ of finite length, and the way that such $\bar{\mathcal{U}}$ fit into $\mathcal{U}$ agree with how $\bar{\mathcal{U}}^{\prime}=i^{\mathcal{T}}\bar{\mathcal{U}}$ fit into $\mathcal{U}^{\prime}=i^{\mathcal{T}}\mathcal{U}$, and $\mathcal{U}^{\prime}$ is also the direct limit of these finite iterations $\bar{\mathcal{U}}^{\prime}$. (Here it is important that we are dealing with the copy $i^{\mathcal{T}}\mathcal{U}$, not the image $i^{\mathcal{T}}(\mathcal{U})$, because if $\mathrm{lh}(\mathcal{U})\geq\theta$ where $\mathrm{cr}(i^{\mathcal{T}})=\theta$, then these things don’t work quite like this for $i^{\mathcal{T}}(\mathcal{U})$.) ∎
2.11 Lemma.
Assume ZFC and let $\kappa$ be a limit of measurable cardinals. Let $\xi\in\mathrm{OR}$. Then there is $\mu<\kappa$ such that $i^{\mathcal{T}}(\xi)=\xi$ for all linear iterations of measures $\mathcal{T}$ on $V$ of successor length such that $\mathcal{T}$ is above $\mu$, $\mathcal{T}$ is based on $V_{\kappa}$ and $i^{\mathcal{T}}(\kappa)=\kappa$.
Proof.
Let us first prove the lemma restricted to trees which are also of length $<\kappa$ and $\mathcal{T}$ is based on $V_{\delta}$ for some $\delta<\kappa$. Suppose this version fails. Then we can pick a sequence $\mu_{n},\delta_{n},\mathcal{T}_{n}\right>_{n<\omega}$ such that:
  1. $\mu_{n}$ is inaccessible and $\mu_{n}<\delta_{n}<\mu_{n+1}$,
  2. $\mathcal{T}_{n}$ is a linear iteration of measures on $V$, is above $\mu_{n}$ and based on $V_{\delta_{n}}$, and has length $<\mu_{n+1}$, and
  3. $i^{\mathcal{T}_{n}}(\xi)>\xi$.
Let $\mathcal{T}_{1}^{\prime}=i^{\mathcal{T}_{0}}\mathcal{T}_{1}$, let $\mathcal{T}_{2}^{\prime}=i^{\mathcal{T}_{1}^{\prime}}\circ i^{\mathcal{T}_{0}}\mathcal{T}_{2}$, etc, defining $\mathcal{T}_{n}^{\prime}$ for all $n<\omega$. Note that $\mathcal{U}=\mathcal{T}_{0}\ \widehat{\ }\ \mathcal{T}_{1}^{\prime}\ \widehat{\ }\ \mathcal{T}_{2}^{\prime}\ \widehat{\ }\ \ldots$ is a linear iteration on $V$. But note that by Lemma 2.10, $i^{\mathcal{T}_{n}^{\prime}}(\xi)=i^{\mathcal{T}_{n}}(\xi)>\xi$. But then $M^{\mathcal{U}}_{\infty}$ is illfounded, a contradiction.
For ordinals $\xi$, let $\mu_{\xi}$ be the least witness for this weak version of the lemma with respect to $\xi$. We claim that $\mu_{\xi}$ also witnesses the full lemma. (We will not prove this by induction; we will just use the fact that the weak version of the lemma holds at all ordinals $\leq\xi$.) So let $\mathcal{T}$ be an iteration of the form considered for the lemma, which is above $\mu_{\xi}$; we want to see $i^{\mathcal{T}}(\xi)=\xi$. We may assume that $\mathcal{T}$ is not based on $V_{\delta}$ for any $\delta<\kappa$, and we have $i^{\mathcal{T}}(\kappa)=\kappa$, and so $i^{\mathcal{T}}``\kappa$ is cofinal in $\kappa$. This easily implies that $\mathcal{T}$ has length $\lambda+1$ for some limit $\lambda$, and for each $\alpha+1<\mathrm{lh}(\mathcal{T})$, $\mathcal{T}\!\upharpoonright\!(\alpha+1)$ is based on $V_{\delta_{\alpha}}$ for some $\delta_{\alpha}<\kappa$. So $i^{\mathcal{T}}_{0\alpha}(\xi)=\xi$ for each such $\alpha$. So fix such an $\alpha$ and let $\eta<\xi$. Let $\mu=(\mu_{\eta})^{M^{\mathcal{T}}_{\alpha}}$ (that is, $\mu$ is what $M^{\mathcal{T}}_{\alpha}$ thinks is the least witness to the weak version of the lemma with respect to $\eta$). Then we can fix $\beta+1\in(\alpha+1,\mathrm{lh}(\mathcal{T}))$ such that $\mathcal{T}\!\upharpoonright\![\beta,\lambda+1)$ is above $i^{\mathcal{T}}_{\alpha\beta}(\mu)$. Now we claim that for all $\gamma\in[\beta,\lambda)$, we have $i^{\mathcal{T}}_{\beta\gamma}(i^{\mathcal{T}}_{\alpha\beta}(\eta))=i^{\mathcal{T}}_{\alpha\beta}(\eta)$. If for a given $\gamma$, we had $\mathcal{T}\!\upharpoonright\![\beta,\gamma+1)\in M^{\mathcal{T}}_{\beta}$, this would be immediate, but in case $\mathcal{T}\!\upharpoonright\![\beta,\gamma+1)\notin M^{\mathcal{T}}_{\beta}$, we need a little more argument: note that $\mathcal{T}\!\upharpoonright\![\beta,\gamma+1)$ can anyway be absorbed into some tree $\mathcal{T}^{\prime}\in M^{\mathcal{T}}_{\beta}$, which is above $i^{\mathcal{T}}_{\alpha\beta}(\mu)$, and which is based on some $V_{\delta}^{M^{\mathcal{T}}_{\beta}}$ with $\delta<\kappa$, and has length $<\kappa$, so $i^{\mathcal{T}^{\prime}}(i^{\mathcal{T}}_{\alpha\beta}(\eta))=i^{\mathcal{T}}_{\alpha\beta}(\eta)$, but therefore $i^{\mathcal{T}}_{\beta\gamma}(i^{\mathcal{T}}_{\alpha\beta}(\eta))=i^{\mathcal{T}}_{\alpha\beta}(\eta)$. (The tree $\mathcal{T}^{\prime}$ is a typical instance of a “universal” iteration; we just use every measure in $M^{\mathcal{T}}_{\beta}$ and its images, each enough times, one after the other, to ensure that we have used some “copy” of $E^{\mathcal{T}}_{\varepsilon}$ for each $\varepsilon\in[\beta,\gamma)$.)
So each $\zeta<i^{\mathcal{T}}(\xi)$ is of the form $\zeta=i^{\mathcal{T}}_{\beta\lambda}(\eta)$ for some $\eta<\xi$ such that $i^{\mathcal{T}}_{\beta\gamma}(\eta)=\eta$ for all $\gamma\in[\beta,\lambda)$. But then just note that $\zeta=i^{\mathcal{T}}_{\beta\lambda}(\eta)=\eta$ for all such $\zeta,\eta$, and so $i^{\mathcal{T}}(\xi)=\xi$. ∎
3 Wellorders above measurable cardinals
3.1 Above a single measurable
In [11], assuming $V=L[U]$ where $U$ is a normal measure on a measurable cardinal $\mu$, Lücke and Schlicht characterized those cardinals $\kappa$ such that there is a good $\Sigma_{1}(\{\kappa\})$ wellorder of $\mathcal{P}(\kappa)$: it is precisely $\kappa$ such that $\kappa\leq\mu^{+}$ or $\kappa$ is non-$\mu$-steady. (They also analyzed the analogous question in the canonical model $L[U_{0},U_{1}]$ for two measurables.) In [11, Question 2], they asked, still assuming $V=L[U]$, given an uncountable cardinal $\kappa$, whether there is a (not necessarily good) $\Sigma_{1}(\{\kappa\})$ wellorder of $\mathcal{P}(\kappa)$. (Of course, their result just mentioned already answered this question positively in the case that $\kappa\leq\mu^{+}$ or $\kappa$ is non-$\mu$-steady.)
The question was answered by [9, Theorem 7.1], where Lücke and Müller prove the following facts, assuming ZFC, $\mu$ is a measurable cardinal and $\kappa>\mu^{+}$ is $\mu$-steady:
  1. (i)
    There is no $\Sigma_{1}(V_{\mu}\cup\{\nu,\nu^{+}\})$ wellorder of $\mathcal{P}(\kappa)$.
  2. (ii)
    If $\operatorname{cof}(\kappa)>\omega$ then there is no $\Sigma_{1}(V_{\mu}\cup\{\kappa\})$ injection from $\kappa^{+}$ into $\mathcal{P}(\kappa)$. [^4] One can also allow the parameter $\kappa^{+}$ without affecting their proof; but if $\nu<\nu^{+}=\kappa$, allowing the parameter $\nu$ would be a problem for their proof.
We show in the following theorem that the hypothesis “if $\operatorname{cof}(\kappa)>\omega$” is not necessary in part (ii) of the Lücke-Müller result. We will also give a slightly different (and more self-contained) proof of part (i), which also yields the slightly stronger fact that there is no such wellorder of $[\kappa]^{\omega}$, nor in fact of $[\eta]^{\omega}_{\uparrow}$, for a certain ordinal $\eta<\kappa$:
3.1 Definition.
For an ordinal $\eta$, $[\eta]^{\omega}_{\uparrow}$ denotes $\{A\in\mathcal{P}(\eta)\bigm{|}A\text{ has ordertype }\omega\}$).
3.2 Theorem.
Assume ZFC and let $\mu$ be a measurable cardinal and $\kappa>\mu^{+}$ be a $\mu$-steady cardinal. Let $\nu>\mu$ be $\mu$-closed and such that $\operatorname{cof}(\nu)\neq\mu$ and $\kappa\in\{\nu,\nu^{+}\}$. Let $p\in V_{\mu}$ and $\varphi$ be a $\Sigma_{1}$ formula. Let $U$ be a measure on $\mu$ and $\eta=\mathrm{cr}(U_{\omega})=i^{\mathcal{T}_{U}}_{0\omega}(\mu)$ (as in 2.1), so $\eta$ is the sup of the “critical sequence” associated to $U$. For $\delta\in\mathrm{OR}$ let $S_{\delta}$ be the class of $U$-$[0,\delta)$-stable sets; note that $V_{\mu}\cup\{\kappa,\kappa^{+}\}\subseteq S_{\kappa}$ and $V_{\mu}\cup\{\nu,\nu^{+},\kappa,\kappa^{+},\eta\}\subseteq S_{\omega}$. Then:
  1. 1.
    There is no $\Sigma_{1}(S_{\omega})$ wellorder of $[\eta]^{\omega}_{\uparrow}$ (and hence no such wellorder of $\mathcal{P}(\kappa)$).
  2. 2.
    There is no $\Sigma_{1}(S_{\kappa})$ injective function $f:\kappa^{+}\to\mathcal{P}(\kappa)$; in fact, for any $\Sigma_{1}(S_{\kappa})$ set $f\subseteq\kappa^{+}\times\mathcal{P}(\kappa)$ and any club $C\subseteq\kappa^{+}$, $f\!\upharpoonright\!C$ is not an injective function (note there is no definability requirement on $C$).
  3. 3.
    Every $\Delta_{1}(S_{\kappa})$ subset of $\kappa$ is in $M^{\mathcal{T}_{U}}_{\kappa}$ (see 2.1); moreover, the supremum of lengths of all $\Sigma_{1}(V_{\mu}\cup\{\kappa,\kappa^{+}\})$ wellorders of $\kappa$ is $<\kappa^{+M^{\mathcal{T}_{U}}_{\kappa}}$.
  4. 4.
    Every $\Delta_{1}(S_{\kappa})$ class $C\subseteq\mathrm{OR}$ is definable from parameters over $M^{\mathcal{T}_{U}}_{\kappa}$, and in particular, $C\cap\alpha\in M^{\mathcal{T}_{U}}_{\kappa}$ for each $\alpha\in\mathrm{OR}$. Moreover, if $C$ is $\Delta_{1}(\{\vec{p}\})$ where $\vec{p}\in(S_{\kappa})^{<\omega}$ then for each ordinal $\beta$, $C\cap\beta$ is $\Delta_{1}^{M^{\mathcal{T}_{U}}_{\kappa}}(\{\vec{q},U_{\kappa},\mathscr{F}\})$ where $(\vec{q},U_{\kappa})=i^{\mathcal{T}_{U}}_{0\kappa}(\vec{p},U)$ and $\mathscr{F}=(^{\kappa}\beta)\cap M^{\mathcal{T}_{U}}_{\kappa}$.
Proof.
Part 1: Let $M_{\alpha}=M^{\mathcal{T}_{U}}_{\alpha}$, $i_{\alpha\beta}=i^{\mathcal{T}_{U}}_{\alpha\beta}$, and $\mu_{\alpha}=i_{0\alpha}(\mu)$. Suppose $<^{*}$ is a wellorder of $[\eta]^{\omega}_{\uparrow}$, $\varphi$ is $\Sigma_{1}$, $\vec{p}\in(S_{\omega})^{<\omega}$ and
$\forall X,Y\in[\eta]^{\omega}_{\uparrow}\ \Big{[}X<^{*}Y\iff\varphi(\vec{p},X,Y)\Big{]}.$ (3)
Write $<^{*}_{n}=i_{0n}({<^{*}})$. Since $i_{0n}(\vec{p},\eta)=(\vec{p},\eta)$ for all $n<\omega$, $M_{n}\models$${<^{*}_{n}}$ is a wellorder of $[\eta]^{\omega}_{\uparrow}$ and line (3) holds”.
Since each $M_{n}$ is closed under $\mu$-sequences, we have $[\eta]^{\omega}_{\uparrow}\cap M_{n}=[\eta]^{\omega}_{\uparrow}$, and by the upward absoluteness of $\Sigma_{1}$, it follows that ${<}^{*}_{n}={<^{*}}$. So from now we just write “${<^{*}}$” instead of “${<}^{*}_{n}$”.
Let $X_{\mathrm{even}}=\{\mu_{2n}\}_{n<\omega}$ and $X_{\mathrm{odd}}=\{\mu_{2n+1}\}_{n<\omega}$. Let $X_{0}=X_{\mathrm{even}}$. Let $X_{1}=\{\mu_{0}\}\cup X_{\mathrm{odd}}$. Let $X_{2}=\{\mu_{0},\mu_{1}\}\cup X_{\mathrm{even}}$. In general let $X_{2k}=\{\mu_{0},\ldots,\mu_{2k-1}\}\cup X_{\mathrm{even}}$ and $X_{2k+1}=\{\mu_{0},\ldots,\mu_{2k}\}\cup X_{\mathrm{odd}}$. Now $\mathrm{cr}(i_{n,n+1})=\mu_{n}$ and $i_{n,n+1}(\mu_{k})=\mu_{k+1}$ for each $k\geq n$. Note that $X_{i}\in M_{n}$ for each $i,n<\omega$, and that $i_{n+1,n+2}(X_{n+1})=X_{n}$.
Since $X_{n+1}\neq X_{n}$, either $X_{n+1}<^{*}X_{n}$ or $X_{n}<^{*}X_{n+1}$. But since $<^{*}$ is wellfounded, there must be some $n$ such that $X_{n}<^{*}X_{n+1}$. Fix such an $n$. Then letting $k_{n+1}=i_{n+1,n+2}$, since $k_{n+1}(X_{n+1})=X_{n}$, we have
$M_{n+1}\models\text{``}k_{n+1}(X_{n+1})<^{*}X_{n+1}\text{''}.$ ” .
Applying $k_{n+1}$ to this statement gives (since $k_{n+1}(k_{n+1})=k_{n+2}$) that
$M_{n+2}\models\text{``}k_{n+2}(k_{n+1}(X_{n+1}))<^{*}k_{n+1}(X_{n+1})\text{''}.$ ” .
Now applying $k_{n+2}$ to the latter statement yields (since $k_{n+2}(k_{n+2})=k_{n+3}$)
$M_{n+3}\models\text{``}k_{n+3}(k_{n+2}(k_{n+1}(X_{n+1})))<^{*}k_{n+2}(k_{n+1}(X_{n+1}))\text{''},$ ” ,
and so on. This produces an infinite descending sequence given by the sequence of images of $X_{n+1}$ under these maps, a contradiction.
Part 2: Let us first consider the case that $C=\kappa^{+}$. Suppose some injection $f:\kappa^{+}\to\mathcal{P}(\kappa)$ is $\Sigma_{1}(S_{\kappa})$. As in [9], it follows that $f\in M^{\mathcal{T}_{U}}_{\alpha}$ for each $\alpha<\kappa$. So as in [9], if $\operatorname{cof}(\kappa)>\omega$ then $M^{\mathcal{T}_{U}}_{\kappa}=\bigcap_{\alpha<\kappa}M^{\mathcal{T}_{U}}_{\alpha}$, so $f\in M^{\mathcal{T}_{U}}_{\kappa}$, which is impossible since $\mathcal{P}(\kappa)\cap M^{\mathcal{T}_{U}}_{\kappa}$ has cardinality $\kappa$. But now we want to generalize this to allow the possibility that $\operatorname{cof}(\kappa)=\omega$.
Well, given $\xi<\kappa^{+}$, we can fix $\alpha<\kappa$ such that $\xi$ is $[\alpha,\kappa)$-stable. Then note that $i^{\mathcal{T}_{U}}_{\alpha\beta}(f(\xi))=f(\xi)$ for all $\beta\in[\alpha,\kappa)$. Let $f(\xi)^{*}=i^{\mathcal{T}_{U}}_{\alpha\kappa}(f(\xi))$ for (any/all) such $\alpha$. So $f(\xi)^{*}\in M^{\mathcal{T}_{U}}_{\kappa}$. Now note that $f(\xi)^{*}\cap\kappa=f(\xi)$, because if $\gamma<\kappa$ then $\gamma\in f(\xi)$ iff, letting $\alpha\in(\gamma,\kappa)$ with $f(\xi)^{*}=i^{\mathcal{T}_{U}}_{\alpha\kappa}(f(\xi))$, we have $i^{\mathcal{T}_{U}}_{\alpha\kappa}(\gamma)\in f(\xi)^{*}=i^{\mathcal{T}_{U}}_{\alpha\kappa}(f(\xi))$, but $\gamma<\alpha\leq\mathrm{cr}(i^{\mathcal{T}_{U}}_{\alpha\kappa})$, so this is iff $\gamma\in f(\xi)^{*}$, as desired. So $f(\xi)\in M^{\mathcal{T}_{U}}_{\kappa}$ for all $\xi<\kappa^{+}$, which as in [9] is a contradiction.
Now in general fix a club $C\subseteq\kappa^{+}$ and suppose $f\subseteq\kappa^{+}\times\mathcal{P}(\kappa)$ is $\Sigma_{1}(S_{\kappa})$ and $f\!\upharpoonright\!C:C\to\mathcal{P}(\kappa)$ is an injective function. Let $C_{\alpha}=i^{\mathcal{T}_{U}}_{0\alpha}(C)$; this is also club in $\kappa^{+}$. So $C^{\prime}=\bigcap_{\alpha<\kappa}C_{\alpha}$ is also. But then for each $\beta\in C^{\prime}$, we get $f(\beta)\in M^{\mathcal{T}_{U}}_{\kappa}$ as before, which is again a contradiction.
Part 3: To see that each $\Delta_{1}(S_{\kappa})$ subset of $\kappa$ is in $M^{\mathcal{T}_{U}}_{\kappa}$, use the proof of part 2. Now each $\Sigma_{1}(S_{\kappa})$ wellorder of $\kappa$ is in fact $\Delta_{1}(S_{\kappa})$. It follows immediately that each such wellorder has length $<\kappa^{+M^{\mathcal{T}_{U}}_{\kappa}}$. But then the supremum of their lengths is also $<\kappa^{+M^{\mathcal{T}_{U}}_{\kappa}}$, since $\operatorname{cof}(\kappa^{+M^{\mathcal{T}_{U}}_{\kappa}})=\mu^{+}$, and we only have $\mu$-many parameters in the set $V_{\mu}\cup\{\kappa,\kappa^{+}\}$ (note that we are not considering arbitrary $\Sigma_{1}(S_{\kappa})$ wellorders of $\kappa$ here).
Part 4: Let $\varphi,\psi$ be $\Sigma_{1}$ formulas and $\vec{p}\in(S_{\kappa})^{<\omega}$ be such that
$\forall\alpha\in\mathrm{OR}\ \Big{[}\varphi(\vec{p},\alpha)\iff\neg\psi(\vec{p},\alpha)\Big{]}.$
Then by elementarity, for all $\beta<\kappa$, $M^{\mathcal{T}_{U}}_{\beta}$ satisfies the same statement, and so by the upward absoluteness of $\Sigma_{1}$, we have
$\forall\alpha\in\mathrm{OR}\ \Big{[}\varphi(\vec{p},\alpha)\iff M^{\mathcal{T}_{U}}_{\beta}\models\varphi(\vec{p},\alpha)\Big{]}.$
It now follows that
$\forall\alpha\in\mathrm{OR}\ \Big{[}\varphi(\vec{p},\alpha)\iff M^{\mathcal{T}_{U}}_{\kappa}\models\varphi(\vec{q},\alpha^{*})\Big{]}$
where $\vec{q}=i^{\mathcal{T}_{U}}_{0\kappa}(\vec{p})$ (here $\alpha^{*}$ denotes *$*_{\kappa}(\alpha)$). The rest now follows easily from Lemma 2.7. Note that we make use of the parameter $({{}^{\kappa}}\beta)\cap M^{\mathcal{T}_{U}}_{\kappa}$ in order to compute *$*_{\kappa}\!\upharpoonright\!\beta$, via that lemma. ∎
3.2 Above infinitely many measurables
In [9, Corollary 7.4], Lücke and Müller showed assuming ZFC + $\kappa$ is a limit a of measurable cardinals:
  1. (i)
    no well ordering of $\mathcal{P}(\kappa)$ is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa,\kappa^{+}\})$, and
  2. (ii)
    if $\operatorname{cof}(\kappa)>\omega$ then no injection $f:\kappa^{+}\to\mathcal{P}(\kappa)$ is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$. [^5] Again, allowing the parameter $\kappa^{+}$ here does no harm.
In [9, Theorem 1.4], they also show, assuming ZFC + $\kappa$ is a limit of measurables, $\operatorname{cof}(\kappa)=\omega$ and there is a wellorder $W$ of some set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$ such that $W$ is $\Sigma_{1}(\{\kappa\})$, that it follows that there is a $\utilde{\Sigma}^{1}_{3}$ wellordering of the reals.
This led them to pose [9, Question 10.4], which asks whether the following theory is consistent: ZFC + “there is a cardinal $\kappa$ which is a limit of measurable cardinals and a family $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$ and a wellorder $<^{*}$ of $D$ such that $<^{*}$ is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$. And [9, Question 10.3] asks the same question, except with the extra demand that $\kappa$ have cofinality $\omega$. The answers (to both) are “no”, as we show in Theorem 3.3 below.
We also sharpen their result (ii) above, by removing the hypothesis that $\operatorname{cof}(\kappa)>\omega$, and also by allowing arbitrary defining parameters in $\mathcal{H}_{\kappa}\cup\mathrm{OR}$; we will also establish an “injective on a club” version.
Lücke and Müller also showed in [9, Theorem 1.1], assuming ZFC + $\kappa$ is a limit of measurable cardinals, that for every $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$, there is a perfect embedding $\iota:{{}^{\operatorname{cof}(\kappa)}}\kappa\to\mathcal{P}(\kappa)$ with $\mathrm{rg}(\iota)\subseteq D$. We will also generalize this theorem, also by allowing any parameters in $\mathcal{H}_{\kappa}\cup\mathrm{OR}$ to be used in defining $D$.
3.3 Theorem.
Assume ZFC and $\kappa\in\mathrm{OR}$ is a limit of measurable cardinals. Then:
  1. 1.
    For every $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$, there is a perfect embedding $\iota:{{}^{\operatorname{cof}(\kappa)}\kappa}\to\kappa$ with $\mathrm{rg}(\iota)\subseteq D$.
  2. 2.
    There is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ wellorder $<^{*}$ of a set $D\subseteq\mathcal{P}(\kappa)$ with $D$ of cardinality $>\kappa$.
  3. 3.
    There are no $f,C$ such that $f\subseteq\kappa^{+}\times\mathcal{P}(\kappa)$, $f$ is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$, $C$ is club in $\kappa^{+}$ and $f\!\upharpoonright\!C$ is an injection.
3.4 Remark.
This theorem also yields a direct proof of the “(i) implies (ii)” part of [9, Theorem 1.5]; in fact we have direct implication in that direction, instead of just relative consistency.
Proof of Theorem 3.3.
Part 3: Note that this is a corollary of Theorem 3.2 part 2 combined with Lemma 2.11.
Parts 1 and 2: Let $\varphi$ be $\Sigma_{1}$ and $\vec{p}\in(\mathcal{H}_{\kappa}\cup\mathrm{OR})^{<\omega}$ and suppose $D=\{X\in\mathcal{P}(\kappa)\bigm{|}\varphi(\vec{p},X)\}$ has cardinality $>\kappa$. Let $\eta=\operatorname{cof}(\kappa)\leq\kappa$. Fix a strictly increasing sequence $\kappa_{\alpha}\right>_{\alpha<\eta}$ of measurable cardinals cofinal in their supremum $\kappa$, such that:
  1. (i)
    $\eta<\kappa_{0}$ if $\eta<\kappa$,
  2. (ii)
    $\vec{p}\cap V_{\kappa}\in V_{\kappa_{0}}$,
  3. (iii)
    $i^{\mathcal{T}}(\xi)=\xi$ for each ordinal $\xi\in\vec{p}$ whenever $\mathcal{T}$ is a linear iteration of measures on $V$ of length $\leq\kappa+1$, which is above $\kappa_{0}$, based on $V_{\kappa}$, and $i^{\mathcal{T}}(\kappa)=\kappa$.
(Use Lemma 2.11 to see there is such a $\kappa_{0}$.) Fix a sequence $U_{\alpha}\right>_{\alpha<\eta}$ of $\kappa_{\alpha}$-complete measures $U_{\alpha}$ on $\kappa_{\alpha}$.
Note that there is some $X\in D$ such that $X\neq i^{\mathcal{T}_{U_{\alpha}}}_{0\kappa}(X\cap\kappa_{\alpha})$ for all $\alpha<\eta$. Fix such an $X$. By thinning out the sequence of measurables, we may assume that for all $\alpha<\eta$, we have
$i^{\mathcal{T}_{U_{\alpha}}}_{0\kappa_{\alpha+1}}(X\cap\kappa_{\alpha})\neq X\cap\kappa_{\alpha+1}.$
Let $A\subseteq\{\kappa_{\alpha}\}_{\alpha<\eta}$ be cofinal in $\kappa$. Write $A=\{\kappa^{A}_{\alpha}\}_{\alpha<\eta}$ with $\kappa^{A}_{\alpha}<\kappa^{A}_{\beta}$ for $\alpha<\beta$. Then define the iteration $\mathcal{U}_{A}$ to be that iteration $\mathcal{U}$ using measures from $\{U_{\alpha}\}_{\alpha<\eta}$ and its pointwise images, with $i^{\mathcal{U}}(\kappa_{\alpha})=\kappa^{A}_{\alpha}$ for each $\alpha<\eta$, and $\mathrm{cr}(E^{\mathcal{U}}_{\alpha})<\mathrm{cr}(E^{\mathcal{U}}_{\beta})$ for $\alpha<\beta<\eta$ (it is easy to see that this uniquely determines $\mathcal{U}$). Then $\mathcal{U}_{A}$ is of the kind mentioned in clause (iii) above (in particular $i^{\mathcal{U}_{A}}(\kappa)=\kappa$), so $i^{\mathcal{U}_{A}}(\vec{p})=\vec{p}$. Let $M_{A}=M^{\mathcal{U}_{A}}_{\infty}$, which is wellfounded; let $i_{A}=i^{\mathcal{U}_{A}}:V\to M_{A}$. Since $\varphi$ is $\Sigma_{1}$, it follows that $i_{A}(X)\in D$.
Now part 1 follows readily from these considerations, as in the proof of [9, Theorem 1.1].
Part 2: We may assume that $<^{*}$ is a $\Sigma_{1}(\{\vec{p}\})$ wellorder of $D$, and $\psi$ is a $\Sigma_{1}$ formula such that for all $X,Y\in\mathcal{P}(\kappa)$, we have $X<^{*}Y\iff\psi(\vec{p},X,Y)$.
Continuing with $A$ as above, suppose that $A\cap\{\kappa_{n}\}_{n<\omega}\subsetneq\{\kappa_{n}\}_{n<\omega}$ but $A\cap\{\kappa_{n}\}_{n<\omega}$ is infinite. Letting $n$ be least such that $\kappa_{n}\notin A$, so $\mathrm{cr}(i_{A})=\kappa_{n}$ and $\kappa_{n+1}\leq i_{A}(\kappa_{n})$, note that $X\neq i_{A}(X)$, and in fact $X\cap\kappa_{n+1}\neq i_{A}(X)\cap\kappa_{n+1}$. A slight generalization of this also shows that if $A^{\prime}$ is likewise and $A\cap\{\kappa_{n}\}_{n<\omega}\neq A^{\prime}\cap\{\kappa_{n}\}_{n<\omega}$ then $i_{A}(X)\neq i_{A^{\prime}}(X)$.
Now we claim that if $A\cap\{\kappa_{n}\}_{n<\omega}\subsetneq\{\kappa_{n}\}_{n<\omega}$ and $A\cap\{\kappa_{n}\}_{n<\omega}$ is infinite then $X<^{*}i_{A}(X)$. For otherwise $i_{A}(X)<^{*}X$, so $\psi(\vec{p},Y,X)$ where $Y=i_{A}(X)$. We can apply $i_{A}$ to this statement, which since $i_{A}(\vec{p})=\vec{p}$, gives
$M_{A}\models\psi(\vec{p},Y^{\prime},X^{\prime})$
where $Y^{\prime}=i_{A}(Y)=i_{A}(i_{A}(X))$ and $X^{\prime}=i_{A}(X)$. So in fact $\psi(\vec{p},Y^{\prime},X^{\prime})$ holds, so $i_{A}(i_{A}(X))<^{*}i_{A}(X)$. But then we can also apply $i_{A}$ to the latter statement, and so on, giving a desceding $\omega$-sequence through $<^{*}$, a contradiction.
It similarly follows that if $A,B$ are as above and $\{\kappa_{\alpha}\}_{\omega\leq\alpha<\eta}\subseteq A\cap B$ and $B\cap\{\kappa_{n}\}_{n<\omega}\subsetneq A\cap\{\kappa_{n}\}_{n<\omega}$, then $i_{A}(X)<^{*}i_{B}(X)$. (We have $B\cap\{\kappa_{n}\}_{n<\omega}\in M_{A}$, and working in $M_{A}$, we can define $M_{B}$ and the factor map $i_{AB}:M_{A}\to M_{B}$, and get $i_{AB}(i_{A}(X))=i_{B}(X)$, from which we get $i_{A}(X)<^{*}i_{B}(X)$ much as before. Here is some more detail, assuming $\eta=\omega$ to ignore irrelevant details. We have $B\in M_{A}$, and $B\subseteq i_{A}(\{\kappa_{n}\}_{n<\omega})=\{\kappa^{A}_{n}\}_{n<\omega}$, and so we can define $(M_{B})^{M_{A}}$ and $(i_{B})^{M_{A}}$ in $M_{A}$; that is, we apply the preceding definitions in $M_{A}$, working from the parameter $i_{A}(\{\kappa_{n}\}_{n<\omega})=\{\kappa^{A}_{n}\}_{n<\omega}$, to define $(M_{B})^{M_{A}}$ and $(i_{B})^{M_{A}}$. A simple instance of normalization [^6] See for example [20], [26]; but here things are much simpler. and yields that $(M_{B})^{M_{A}}=M_{B}$ and $i_{B}=(i_{B})^{M_{A}}\circ i_{A}$. Writing $i_{AB}=(i_{B})^{M_{A}}$, we then have $i_{AB}:M_{A}\to M_{B}$ and $i_{AB}(i_{A}(X))=i_{B}(X)$.)
But now let $B=\{\kappa_{2n}\}_{n<\omega}\cup\{\kappa_{\alpha}\}_{\omega\leq\alpha<\eta}$, and consider $A_{n}=B\cup\{\kappa_{1},\ldots,\kappa_{2n-1}\}$. Then $A_{n}\subsetneq A_{n+1}$ for all $n$, so $i_{A_{n+1}}(X)<^{*}i_{A_{n}}(X)$ for all $n$, giving a strictly descending $\omega$-sequence through $<^{*}$, a contradiction. ∎
3.3 Above rank-to-rank
3.5 Theorem.
Assume ZFC + $I_{2}$ holds at $\kappa$. Then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\mathcal{H}_{\kappa}\}\cup\mathrm{OR})$ wellorder of a set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$.
Note that the crucial difference between this and the results on measurable cardinals is that we also have $\mathcal{H}_{\kappa}$ itself available as a parameter, not just elements of $\mathcal{H}_{\kappa}$.
Proof.
Let $\vec{p}\in\mathcal{H}_{\kappa}\cup\{\mathcal{H}_{\kappa}\}\cup\mathrm{OR}$ and $\varphi$ be $\Sigma_{1}$, and suppose that $\varphi(\vec{p},\cdot,\cdot)$ defines a wellorder of a set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$. Let $j:V\to M$ witness $I_{2}$ with $j(\vec{p})=\vec{p}$. (Let $k:V\to M$ witness $I_{2}$ at $\kappa$, and be given by a rank-to-rank extender over $V_{\kappa}$. Let $k_{0}=k$ and $k_{n+1}=k_{n}(k_{n})$. Then for each $\alpha\in\mathrm{OR}$, there is $n<\omega$ such that $k_{n}(\alpha)=\alpha$.)
Let $j_{n+1}=j_{n}(j_{n})$. Let $\mu_{n}=\mathrm{cr}(j_{n})$. Define sets $X_{n}$ as in the proof of Theorem 3.2 part 1, and now continue as in that proof to a contradiction; the key point is that $j(\mathcal{H}_{\kappa})=\mathcal{H}_{\kappa}$, so $j_{n}(\mathcal{H}_{\kappa})=\mathcal{H}_{\kappa}$ for all $n<\omega$. ∎
3.6 Remark.
Similar methods give that if $I_{2}$ holds at $\kappa$ and $D$ is $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\mathcal{H}_{\kappa}\}\cup\mathrm{OR})$ and has cardinality $>\kappa$, then $D$ has a perfect subset, appropriately defined.
4 Descriptive properties of $L[U]$
Throughout this section, we assume ZFC, $\mu$ is a measurable cardinal and $V=L[U]$ where $U$ is a normal measure on $\mu$. We make some observations on descriptive set theory at uncountable cardinals $\kappa$ in this context, using the basic theory of $L[U]$ (see [7], [6]), by which, $\mu$ is the unique measurable and $U$ the unique normal measure on $\mu$. Note that since GCH holds in $L[U]$, a cardinal $\kappa$ is $\mu$-steady iff $\kappa>\mu^{+}$, $\operatorname{cof}(\kappa)\neq\mu$ and $\kappa$ is not of the form $\nu^{+}$ where $\nu$ is a cardinal with $\operatorname{cof}(\nu)=\mu$. Say a filter $D$ is $\kappa$-good iff $D$ is a filter over some $\gamma<\kappa$ and $D\in L[D]\models$$D$ is a normal measure on $\gamma$”. We officially work with $L[U]$ in the modern, fine structural hierarchy, using notation as described in §1.1. (So the extender sequence $\mathbb{E}^{L[U]}$ of $L[U]$ contains not only $U$, but also many partial measures, on many cardinals $\leq\mu$.)
[…]
Part 4 of the following theorem should be contrasted with Theorem 3.2 part 2, and part 5 already follows from [9, Theorem 1.2] in the case that $\kappa$ is singular (in $L[U]$). But we will give a self-contained proof.
4.1 Theorem.
Assume $V=L[U]$ where $U$ is a normal measure and let $\mu$ be measurable. Then for every uncountable cardinal $\kappa$, we have:
  1. 1.
    There is a set $X\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ and a wellorder $<^{*}$ of $X$ such that $X,{<^{*}}$ are both $\Sigma_{1}(\{\kappa\})$.
  2. 2.
    By [11], if either $\kappa\leq\mu^{+}$ or $\kappa$ is non-$\mu$-steady then $\mathbb{E}^{L[U]}\!\upharpoonright\!\kappa^{+}$ and $<_{L[U]}\!\upharpoonright\!\mathcal{H}_{\kappa^{+}}$ are $\Sigma_{1}(\{\kappa\})$, so the set of master codes of levels of $L[U]$ (in the fine hierarchy given by $\mathbb{E}^{L[U]}$) projecting to $\kappa$ is $\Sigma_{1}(\{\kappa\})$.
  3. 3.
    If $\kappa\leq\mu^{+}$ or $\kappa$ is non-$\mu$-steady then there is a $\Sigma_{1}(\{\kappa\})$ function $f:\kappa^{+}\to\mathcal{P}(\kappa)$ such that for each $\xi<\kappa^{+}$, $f(\xi)$ is a pair $(T,A)$ such that for some $\eta\in(\xi,\kappa^{+})$, we have $A\in(L[U]|\eta)\models$$\kappa$ is the largest cardinal and $A$ is the $<_{L[U]}$-least wellorder of $\kappa$ in ordertype $\xi$”,
    $L[U]|\eta=\mathrm{Hull}^{L[U]|\eta}(\kappa),$ (4)
    $T$ is the theory
    $T=\operatorname{Th}^{L[U]|\eta}(\kappa),$ (5)
    and if $\kappa$ is singular then $L[U]|\eta\models$$\kappa$ is singular”. [^7] Here when defining the hull $\mathrm{Hull}^{L[U]|\eta}(\kappa)$ and the theory $\operatorname{Th}^{L[U]|\eta}(\kappa)$, the language is that of premice; in particular, there are symbols for $\mathbb{E}^{L[U]|\eta}$ and the active extender $F^{L[U]|\eta}$. We encode finite tuples of ordinals via some standard Gödel coding, so that we can simply talk about tuples $(\varphi,\vec{x})$ where $\varphi$ is a formula and $\vec{x}$ a finite tuple of parameters in $\kappa$. Likewise, $(A,T)$ is encoded simply as a single subset of $\kappa$.
  4. 4.
    If $\kappa>\mu^{+}$ and $\kappa$ is $\mu$-steady then there is a set $d\subseteq\kappa^{+}$ which is stationary-co-stationary in $\kappa^{+}$, and a function $f:d\to\mathcal{P}(\kappa)$, such that $d,f$ are both $\Sigma_{1}(\{\kappa\})$, and for each $\xi\in d$, $f(\xi)$ is a pair $(T,A)$ such that for some $\kappa$-good filter $D$ and some ordinal $\eta\in(\kappa,\kappa^{+})$, we have $A\in(L[D]|\eta)\models$$\kappa$ is the largest cardinal and $A$ is the $<_{L[D]}$-least wellorder of $\kappa$ in ordertype $\xi$”,
    $L[D]|\eta=\mathrm{Hull}^{L[D]|\eta}(\kappa),$ (6)
    $T$ is the theory
    $T=\operatorname{Th}^{L[D]|\eta}(\kappa),$ (7)
    and if $\kappa$ is singular then $L[D]|\eta\models$$\kappa$ is singular”.
  5. 5.
    There is a $\Sigma_{1}(\{\kappa\})$ set $X\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$, such that $X$ has no simply perfect subset.
Proof.
Part 1 will be an immediate corollary of part 4, taking $X=\mathrm{rg}(f)$ and for $A,B\in X$, setting $A<^{*}B$ iff $f^{-1}(A)<f^{-1}(B)$.
Part 2: See [11].
Part 3: This is a routine consequence of part 2 (and is a simplification of part 4, modulo part 2).
Part 4: Define $f$ as the set of pairs $(\xi,(T,A))$ such that $\kappa<\xi<\kappa^{+}$ and there is a $\kappa$-good filter $D$ and an ordinal $\eta\in(\xi,\kappa^{+})$ such that $A\in(L[D]|\eta)\models$$\mathrm{card}(\xi)=\kappa$”, $A$ is the $<_{L[D]}$-least wellordering of $\kappa$ in ordertype $\xi$, $\xi$ is $D$-$1$-stable but non-$D$-00-stable, if $\kappa$ singular then $L_{\eta}[D]\models$$\kappa$ is singular”, there is no $\eta^{\prime}<\eta$ with these properties with respect to $D,\xi$, and $T$ is the theory as in line (5). Because of the minimality of $\eta$, line (6) then also holds.
Claim 1.
There are stationarily many $\xi<\kappa^{+}$ such that there are $T,A$ with $(\xi,(T,A))\in f$, as witnessed by $D=U$ and some $\eta$.
Proof.
By Lemma 2.5, for each $\xi\leq\kappa^{+}$, there is $\alpha<\kappa$ such that $\xi$ is $\alpha$-stable. So for some $\alpha<\kappa$, there are cofinally many $\xi<\kappa^{+}$ which are $\alpha$-stable. Then since $\kappa,\kappa^{+}$ are 00-stable, it follows that in fact there are cofinally many $\xi<\kappa^{+}$ which are 00-stable.
Let $\eta_{0}$ be the least $\eta>\kappa$ such that if $\kappa$ is singular then $L[U]\models$$\kappa$ is singular”. Now it suffices to see that there are stationarily many $\xi\in[\eta_{0},\kappa^{+})$ which are $1$-stable but non-00-stable, since for all such $\xi$, there is $(T,A)$ as claimed. So let $C\subseteq\kappa^{+}$ be club in $\kappa^{+}$. Let $\xi$ be the least ordinal $\geq\eta_{0}$ of cofinality $\mu$ which is a limit point of $C$ (so $\xi\in C$ also) and is a limit of 00-stable ordinals. Because $\operatorname{cof}(\xi)=\mu$, $\xi$ is non-00-stable and $i^{\mathcal{T}_{U}}_{1\beta}$ is continuous at $\xi$ for each $\beta\in[1,\kappa)$ (we have $\operatorname{cof}^{M^{\mathcal{T}_{U}}_{1}}(\xi)=\operatorname{cof}(\xi)=\mu$, so $\operatorname{cof}^{M^{\mathcal{T}_{U}}_{\beta}}(i^{\mathcal{T}_{U}}_{1\beta}(\xi))=\mu$ for all $\beta\in[1,\kappa)$). But also because $\xi$ is a limit of 00-stable ordinals (hence also $1$-stable), it follows that $\xi$ is $1$-stable, which suffices. ∎
Claim 2.
$f$ is a function.
Proof.
Let $\xi<\kappa^{+}$ and $A,T,A^{\prime},T^{\prime}$ be such that $(\xi,(A,T))\in f$, as witnessed by $D,\eta$, and $(\xi,(A^{\prime},T^{\prime}))\in f$, as witnessed by $D^{\prime},\eta^{\prime}$. Then observe that $D=D^{\prime}$, $A=A^{\prime}$, $\eta=\eta^{\prime}$, and $T=T^{\prime}$, as desired.∎
Let $d=\mathrm{dom}(f)$.
Claim 3.
$d$ is co-stationary in $\kappa^{+}$.
Proof.
By Theorem 3.2 part 2, or more directly, just observe (by a simplification of the proof of Claim 1) that there are stationarily many $\xi<\kappa^{+}$ which are 00-stable.∎
It just remains to verify:
Claim 4.
$d,f$ are $\Sigma_{1}(\{\kappa\})$.
Proof.
It is straightforward to see that most of the defining clauses are $\Sigma_{1}(\{\kappa\})$, but there is a subtlety regarding the clause “$\xi$ is $D$-$1$-stable”. Let $\theta=\operatorname{cof}^{L[D]}(\kappa)$. Note that by the requirements on $(\xi,(T,A)),\eta,D$, $\eta$ is the least $\eta^{\prime}>\xi$ such that $L[D]|\eta^{\prime}\models$$\mathrm{card}(\xi)=\kappa$ and $\operatorname{cof}(\kappa)=\theta$”, and therefore, $\eta^{\prime}=\eta^{\prime\prime}+\omega$ for some $\eta^{\prime\prime}$ such that $L[D]|\eta^{\prime\prime}$ projects to $\kappa$, and there is a bijection $\pi:\kappa\to\xi$ and a cofinal function $h:\theta\to\kappa$ which are both definable from parameters over $L[D]|\eta^{\prime\prime}$. Let $\mu_{D}=\mathrm{cr}(D)$. We have $\mu_{D}<\kappa$, so $\mathcal{P}(\mu_{D})\cap L[D]\subseteq L[D]|\kappa\subseteq L[D]|\eta$.
Now it suffices to see that for all functions $g:\mu_{D}\to\eta$ with $g\in L[D]$, there is $X\in D$ such that $g\!\upharpoonright\!X\in L[D]|\eta$, since then $i^{\mathcal{T}_{D}^{L[D]}}_{0\alpha}(\xi)=i^{\mathcal{T}^{L[D]|\eta}_{D}}_{0\alpha}(\xi)$ for all $\alpha<\kappa$, where $\mathcal{T}^{L[D]|\eta}_{D}$ is the linear iteration of $L[D]|\eta$ (whereas $\mathcal{T}_{D}^{L[D]}$ is the linear iteration of $L[D]$). But because $\kappa$ is $\mu$-steady, we have $\operatorname{cof}(\kappa)\neq\mu$, and it follows that $\theta=\operatorname{cof}^{L[D]}(\kappa)\neq\mu_{D}$. (Indeed, if $D\neq U$ then we can fix $\alpha<\kappa$ such that $(D,\mu_{D})=i^{\mathcal{T}_{U}}_{0\alpha}(U,\mu)$. But $i^{\mathcal{T}_{U}}_{0\alpha}(\kappa)=\kappa$ since $\kappa$ is $\mu$-steady, so letting $\theta^{\prime}=\operatorname{cof}(\kappa)\neq\mu$, we have $i^{\mathcal{T}_{U}}_{0\alpha}(\theta^{\prime})=\theta$, so $\theta\neq\mu_{D}$.)
Now let $g:\mu_{D}\to\eta$ with $g\in L[D]$. Then there is $X_{0}\in D$ such that $g\!\upharpoonright\!X_{0}$ is bounded in $\eta$, since $\eta=\eta^{\prime}+\omega$. Thus, we might as well assume that $g:\mu_{D}\to\xi$. Let $\pi:\kappa\to\xi$ be a bijection in $L[D]|\eta$. Let $g^{\prime}=\pi^{-1}\circ g:\mu_{D}\to\kappa$. Because $\operatorname{cof}^{L[D]}(\kappa)\neq\mu_{D}$, there is $X\in L[D]$ such that $g^{\prime}\!\upharpoonright\!X$ is bounded in $\kappa$, and therefore $g^{\prime}\!\upharpoonright\!X\in L[D]|\kappa$. But then $g\!\upharpoonright\!X=\pi\circ g^{\prime}\!\upharpoonright\!X\in L[D]|\eta$, as desired. ∎
Part 5: We have two cases:
Case 1.
$\mu^{+}<\kappa$ and $\kappa$ is $\mu$-steady.
Let $f:d\to\mathcal{P}(\kappa)$ be as in part 4 and $X=\mathrm{rg}(f)$. We claim this set works. For clearly $X$ is $\Sigma_{1}(\{\kappa\})$ and $\mathrm{card}(X)=\kappa^{+}$, so we just need to see that $X$ has no simply perfect subset.
Suppose not. Let $\theta=\operatorname{cof}(\kappa)$ and $\sigma:{{}^{<\theta}}2\to{{}^{<\kappa}}2$ be a simply perfect function with $(\bigcup_{\alpha<\theta}\sigma(x\!\upharpoonright\!\alpha))\in X$ for all $x\in{{}^{\theta}}2$. Let $\mathbb{P}=\mathrm{Add}(\theta,1)$. Note that if $\theta=\omega$, this is just Cohen forcing. Then $\mathbb{P}$ is $\theta$-closed. Let $g$ be $(V,\mathbb{P})$-generic. Then $({{}^{<\theta}}V)\cap V[g]\subseteq V$.
Work in $V[g]$. Let $x=(\bigcup g)\in{{}^{\theta}}2$. Let $(A,T)=\bigcup_{\alpha<\theta}\sigma(x\!\upharpoonright\!\alpha)$. We have $(A,T)\notin V$, since $x\notin V$, and $(A,T)$ determines $x$ via $\sigma$. But we will observe that $(A,T)$ satisfies the conditions required for elements of $X$ (except that it is in $V[g]\backslash V$); that is, there are some (uniqely determined) corresponding $D,\eta,\xi$ with those properties holding in $V[g]$. But then by the uniqueness of $L[U]$ and its iterates, we get $D,T,A\in V$, and therefore $x\in V$, a contradiction.
[…]
Let $M_{T}$ be the model determined by $T$. Suppose first that $\theta>\omega$. Then $M_{T}$ is wellfounded, since then every countable substructure of $M_{T}$ can be realized in the model $L[D^{\prime}]|\eta^{\prime}$ corresponding to some branch in $V$. So $M_{T}=L[D]|\eta$ where $L[D]|\eta\models$$D$ is a $\kappa$-good filter”. Since $\omega_{1}^{V[G]}=\omega_{1}^{V}<\kappa\subseteq M_{T}$, it also follows that $D$ is truly iterable, and so $(A,T)$ does indeed satisfy the requirements for elements of $X$, yielding the desired contradiction.
Now suppose that $\theta=\omega$. We use a different argument to see that $M_{T}$ is wellfounded. Working in $V$, consider the tree $S$ of attempts to build a real $y\in{{}^{\omega}}2$ and a descending sequence of ordinals through the model $M_{y}$ determined by $\bigcup_{n<\omega}(\sigma(y\!\upharpoonright\!n))$. (Here the $n$th level of $S$ specifies some finite segment of $y$ of length $\geq n$, and terms defining the first $n$ elements of the desired descending sequence through (the to-be-completed) $\mathrm{OR}^{M_{y}}$.) Then since (in $V$) every branch through $\sigma$ gives a wellfounded model, $S$ is wellfounded. It follows that $M_{T}$ is wellfounded. The rest of the argument is as in the case that $\theta>\omega$ (the value of $\theta$ was not relevant to the iterability of $D$; we still have $\omega_{1}^{V[G]}=\omega_{1}^{V}<\kappa\subseteq M_{T}$).
Case 2.
$\omega<\kappa\leq\mu^{+}$ or [$\mu^{+}<\kappa$ and $\kappa$ is non-$\mu$-steady].
This is a slight variant of the previous case, substituting the calculation of [11] to identify $\mathbb{E}^{L[U]}$ appropriately, instead of the methods we used above. We leave the details to the reader. ∎
4.2 Definition.
Fix an uncountable cardinal $\kappa$. Relative to $\kappa$, and some definability class $\Gamma$, we say that $A\subseteq\kappa$ is a $\Gamma$-singleton iff $\{A\}$ is $\Gamma$-definable. As usual, we say that some class $C$ of singletons forms a basis for some definability class $\Gamma$ iff every non-empty $X\subseteq{{}^{\kappa}}\kappa$ which is $\Gamma$-definable has an element $x\in C$.
4.3 Theorem.
Assume $V=L[U]$ where $U$ is a normal measure on $\mu$. Let $\kappa>\mu^{+}$ be a $\mu$-steady cardinal. Then:
  1. 1.
    For every 00-stable ordinal $\alpha$, every $\Sigma_{1}(\{\alpha\})$-singleton is in $M^{\mathcal{T}_{U}}_{\kappa}$. In particular, every $\Sigma_{1}(\{\kappa\})$-singleton is in $M^{\mathcal{T}_{U}}_{\kappa}$.
  2. 2.
    Not every $\Pi_{1}(\{\kappa\})$ singleton is in $M^{\mathcal{T}_{U}}_{\kappa}$.
  3. 3.
    The $\Sigma_{1}(\{\kappa\})$ singletons do not form a basis for $\Delta_{1}(\{\kappa\})$ (thus, neither $\Sigma_{1}(\{\kappa\})$ nor $\Delta_{1}(\{\kappa\})$ have the uniformization property).
  4. 4.
    In fact, the $\Pi_{1}(\{\kappa\})$ singletons do not form a basis for $\Sigma_{1}(\{\kappa\})$.
Proof.
Part 1: This is an immediate corollary of Theorem 3.2 part 3, since for $A$ a set of ordinals, $A$ is a $\Sigma_{1}(S)$-singleton iff $A$ is $\Delta_{1}(S)$.
Part 2: $U$, coded naturally as a subset of $\kappa$ (using the $L[U]$-order of constructibility) is a such a $\Pi_{1}(\{\kappa\})$ singleton, since it is the unique $\kappa$-good filter $D$ such that every bounded subset of $\kappa$ is in $L_{\kappa}[D]$.
Part 3: Let $X=\mathcal{P}(\kappa)\backslash M^{\mathcal{T}_{U}}_{\kappa}$. Then $X$ is $\Delta_{1}(\{\kappa\})$, since the following three conditions are equivalent: (i) $A\in X$, (ii) $A\subseteq\kappa$ and there is a $\kappa$-good filter $D$ such that $A\notin M^{\mathcal{T}_{D}^{L[D]|\kappa}}_{\kappa}$, and (iii) $A\subseteq\kappa$ and for every $\kappa$-good filter $D$, we have $A\notin M^{\mathcal{T}_{D}^{L[D]|\kappa}}_{\kappa}$. (Here $\mathcal{T}^{L[D]|\kappa}_{D}$ is the iteration on $L[D]|\kappa$ given by iterating $D$.) But by part 1, there is no $\Sigma_{1}(\{\kappa\})$ singleton $A\in X$.
Part 4: Let $X$ be the set of all $A\subseteq\kappa$ such that for some limit ordinal $\eta<\kappa$, we have $A\in\bigcap_{\alpha<\eta}M^{\mathcal{T}_{U}}_{\alpha}$ but $A\notin M^{\mathcal{T}_{U}}_{\eta}$. Note that by Lemma 2.8, $X$ is $\Sigma_{1}(\{\kappa\})$. Suppose there is a $\Pi_{1}(\{\kappa\})$ singleton $A\in X$, and let $\psi$ be a $\Pi_{1}$ formula such that $A$ is the unique $B\subseteq\kappa$ such that $\psi(\kappa,B)$. Let $\eta$ be least such that $A\notin M^{\mathcal{T}_{U}}_{\kappa}$; so $\eta$ is a limit, as $A\in X$. By $\Pi_{1}$ downward absoluteness, for all $\alpha<\eta$, $M^{\mathcal{T}_{U}}_{\alpha}\models\psi(\kappa,A)$. But by elementarity of $i^{\mathcal{T}_{U}}_{0\alpha}$, therefore $M^{\mathcal{T}_{U}}_{\alpha}\models$$A$ is the unique $B\subseteq\kappa$ such that $\psi(\kappa,B)$”. It follows that $i^{\mathcal{T}_{U}}_{0\alpha}(A)=A$. So $A$ is $[0,\eta)$-stable, so by Lemma 2.8, $A\in M^{\mathcal{T}_{U}}_{\eta}$, a contradiction. ∎
5 Simply definable good wellorders in 1-small mice
For mice $M$ and $M$-cardinals $\kappa$, the standard $M$-order of constructibility $<^{M}$ restricts to a good wellorder of $\mathcal{H}^{M}_{\kappa^{+M}}=\left\lfloor M|\kappa^{+M}\right\rfloor$ (see [11] for the definition of good, for example). In [11], Lücke and Schlicht characterized the uncountable cardinals $\kappa$ of $L[U]$ for which there is a $\Sigma_{1}(\{\kappa\})$ good wellorder of $\mathcal{H}^{L[U]}_{\kappa^{+L[U]}}$. In the cases that these exist, the restriction of $<^{L[U]}$ is itself good and $\Sigma_{1}(\{\kappa\})$. Since under ZFC, if $\kappa$ is a limit of measurable cardinals then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ wellorder of $\mathcal{P}(\kappa)$, the following result is close to optimal:
5.1 Theorem.
Let $M$ be a 1-small mouse. Let $\kappa$ be an uncountable cardinal of $M$ which is not a limit of measurable cardinals of $M$ and such that $M\models$$\kappa^{+}$ exists” (that is, $\kappa^{+M}<\mathrm{OR}^{M}$). Suppose that for every $N\triangleleft M|\kappa^{+M}$ and every $\mathcal{T}\in M$ which is a limit length $\omega$-maximal tree on $N$ (so $\delta(\mathcal{T})<\mathrm{OR}^{M}$), letting $\eta$ be such that $\delta(\mathcal{T})+\eta=\mathrm{OR}^{M}$, it is not the case that $\mathcal{J}_{\eta}[M(\mathcal{T})]\models$$\delta(\mathcal{T})$ is Woodin”. Then there is a good wellorder of $\mathcal{H}_{\kappa^{+M}}^{M}$ which is $\Sigma_{1}^{\left\lfloor M\right\rfloor}(\mathcal{H}_{\kappa}^{M}\cup\{\kappa\})$.
It is of note that in case $\kappa$ is a limit cardinal of $M$, the (good) wellorder we use to prove the theorem is not just the restriction $<^{*}$ of the standard order of constructibility $<^{M}$ of $M$ to $\mathcal{H}^{M}_{\kappa^{+M}}$; we do not know whether $<^{*}$ is also so definable. Instead it will involve both the mouse order and orders of constructibility.
For the proof we use the following definition, which is based on standard arguments for absoluteness of iterability in the absence of a proper class model with a Woodin cardinal, like those in [23, §2], and like [1, Lemma 2.1]. Its role here will be analogous to that in Väänänen-Welch [27]:
5.2 Definition.
Let $N$ be a sound $1$-small premouse and $\xi$ be an ordinal such that $L_{\xi}[N]\models$$\omega_{1}$ exists”. Working in $L_{\xi}[N]$, let $S$ be the tree of attempts to build, via finite approximations, a tuple
$(\bar{N},\sigma,\mathcal{T},\rho,\left<\rho_{\alpha}\right>_{\alpha+1<\mathrm{lh}(\mathcal{T})},\iota,Q,\pi)$
such that:
  1. 1.
    $\bar{N}$ is a structure in the language of premice, whose universe is $\omega$,
  2. 2.
    $\sigma:\bar{N}\to N$ is elementary,
  3. 3.
    $\mathcal{T}$ is a putative $\omega$-maximal iteration tree on $\bar{N}$,
  4. 4.
    $\rho:\mathrm{lh}(\mathcal{T})\to\omega_{1}^{L_{\xi}[N]}$ is order-preserving,
  5. 5.
    $\rho_{\alpha}:\mathrm{OR}(M^{\mathcal{T}}_{\alpha})\to\omega_{1}^{L_{\xi}[N]}$ is order-preserving, for each $\alpha+1<\mathrm{lh}(\mathcal{T})$,
  6. 6.
    if $\mathrm{lh}(\mathcal{T})=\alpha+1$ then:
    1. $\iota$ is a strictly descending sequence through $\mathrm{OR}(M^{\mathcal{T}}_{\alpha})$,
    2. if $\alpha$ is a limit, then $Q\triangleleft M^{\mathcal{T}}_{\alpha}$ is a Q-structure for $M(\mathcal{T}\!\upharpoonright\!\alpha)$ and $\pi:\mathrm{OR}^{Q}\to\omega_{1}^{L_{\xi}[N]}$ is order-preserving,
  7. 7.
    if $\mathrm{lh}(\mathcal{T})$ is a limit,
    1. (a)
      $Q=\mathcal{J}_{\gamma}(M(\mathcal{T}))$ for some ordinal $\gamma$, and either $\rho_{n+1}^{Q}<\delta(\mathcal{T})\leq\rho_{n}^{Q}$ for some $n<\omega$, or there is a failure of Woodinness of $\delta(\mathcal{T})$ definable over $Q$,
    2. (b)
      $\pi:\mathrm{OR}^{Q}\to\omega_{1}^{L_{\xi}[N]}$ is order-preserving,
    3. (c)
      $\iota$ is a ranking into $\omega_{1}^{L_{\xi}[N]}$ of the tree of attempts to build a $\mathcal{T}$-cofinal branch $c$ with $Q\trianglelefteq M^{\mathcal{T}}_{c}$. or $M^{\mathcal{T}}_{c}\trianglelefteq Q$.
Note that we can form $S$ as a tree on $\omega\times\max(\mathrm{OR}^{N},\omega_{1}^{L_{\xi}[N]})$.
We say that $N$ is strongly iterable in $L_{\xi}[N]$ iff $L_{\xi}[N]\models$“there is a ranking of $S$ into the ordinals”.
Proof of Theorem 5.1.
The case that $\kappa$ is a successor cardinal is easier, and we deal with that first.
Case 1.
$\kappa=\gamma^{+M}$ for some $M$-cardinal $\gamma\geq\omega$.
In this case, the restriction of the usual order of constructibility will be $(\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\}))^{M}$.
Subcase 1.1.
There is a cutpoint $\eta$ of $M$ such that $\gamma\leq\eta<\kappa=\gamma^{+M}$. (Note this includes the case that $\omega=\gamma=\eta<\kappa=\omega_{1}^{M}$.)
Fix such a cutpoint $\eta$. Let $x=M|\eta$. Then $M|\kappa$ is just the stack of sound premice $N$ such that $M|\eta\triangleleft N$, $\rho_{\omega}^{N}\leq\gamma$, $\eta$ is a cutpoint of $N$ and there is an ordinal $\xi>\kappa$ such that $L_{\xi}[N]\models$$N$ is strongly iterable” (note that then $\omega_{1}^{M}\in L_{\xi}[N]$ and $\omega_{1}^{L_{\xi}[N]}=\omega_{1}^{M}$). This yields a definition of $N|\kappa=N|\gamma^{+M}$ which is $\Sigma_{1}(\{x,\kappa\})$.
We now define $M|\kappa^{+M}$ (as a premouse). Since $\kappa$ is regular in $M$, $M|\kappa^{+M}$ is just the Jensen stack over $M|\kappa$, as computed in $M$ [^8] This is the stack of all sound premice $P$ such that $M|\kappa\triangleleft P$, $\rho_{\omega}^{P}=\kappa$, and $P$ satisfies condensation. Cf. [19]. This is also $\Sigma_{1}(\{x,\kappa\})$, and yields a $\Sigma_{1}^{\left\lfloor M\right\rfloor}(\{x,\kappa\})$ definition of $M|\kappa^{+M}$, and hence of the usual order of constructibility over $\mathcal{H}^{M}_{\kappa^{+M}}$. (Note $\kappa$ might not be a cutpoint of $M$; equivalently, $\gamma$ might be measurable in $M$.)
Subcase 1.2.
Otherwise (there is no such cutpoint).
Set $x=M|\gamma$. For all $\eta\in[\gamma,\kappa=\gamma^{+M})$, there is an extender $E\in\mathbb{E}^{M}$ with $\mathrm{cr}(E)<\gamma\leq\eta<\mathrm{lh}(E)$, and note that any such $E$ is $M$-total. By [18, Corollary 2.18], [19, Theorem 1.4], it follows that for sound premice $N$ such that $M|\gamma\trianglelefteq N$ and $\rho_{\omega}^{N}=\gamma$, we have $N\triangleleft M|\kappa$ iff there is an $M$-total extender $E\in M$ (not required to be in $\mathbb{E}^{M}$) such that $\mu=\mathrm{cr}(E)<\gamma$ and $\mathrm{Ult}(M|\mu^{+M},E)$ is wellfounded and $N\triangleleft\mathrm{Ult}(M|\mu^{+M},E)$. Note that this condition is $\Sigma_{1}(\{x,\kappa\})$, and this suffices to define $M|\kappa=M|\gamma^{+M}$.
From here we compute $M|\kappa^{+M}$ as before.
Case 2.
$\kappa$ is a limit cardinal of $M$.
Subcase 2.1.
There is a cutpoint $\eta$ of $M$ with $\delta<\eta<\kappa$, where $\delta$ is the supremum of $\omega$ and all measurable cardinals of $M|\kappa$.
Take $\eta$ such, and we may assume that $\eta=\theta^{+M}$ for some $\theta$. Set $x=M|\eta$. Working in $M$, say that a premouse $N$ is good iff $M|\eta\triangleleft N$, $\eta$ is a cutpoint of $N$, $\kappa<\mathrm{OR}^{N}$, $\rho_{\omega}^{N}=\kappa$, $N$ is sound, $N\models$“there are no measurable cardinals in the interval $[\eta,\kappa)$”, there is $\xi<\kappa^{+}$ such that $L_{\xi}[N]\models$$N$ is strongly iterable”, and there is $\gamma\in(\kappa,\mathrm{OR}^{N})$ such that $L_{\gamma}[N|\kappa]$ is admissible. Given a good $N$, let $\gamma^{N}$ be the least $\gamma$ such that $L_{\gamma}[N|\kappa]$ is admissible.
The following claim follows from a routine analysis of comparison of $M$ with $N$, which can be executed working inside $M$:
Claim 1.
Let $N$ be good. Then either:
  1. $N\trianglelefteq M$, or
  2. there is an $\omega$-maximal iteration tree $\mathcal{T}$ on $M$ such that:
    1. (a)
      $\mathcal{T}$ is linear,
    2. (b)
      $\mathcal{T}$ is above $\eta$, based on $M|\kappa$, and $\mathrm{lh}(\mathcal{T})=\kappa+1$,
    3. (c)
      $N\triangleleft M^{\mathcal{T}}_{\kappa}$,
    4. (d)
      $\mathcal{T}\in M$,
    5. (e)
      $1\in\mathscr{D}^{\mathcal{T}}$,
    6. (f)
      for every $\alpha+1<\mathrm{lh}(\mathcal{T})$, $\eta<\mathrm{cr}(E^{\mathcal{T}}_{\alpha})$ and $M^{\mathcal{T}}_{\alpha}|\mathrm{cr}(E^{\mathcal{T}}_{\alpha})\models$“There are no measurable cardinals $\geq\eta$”,
    7. (g)
      for all $\alpha+1\leq\beta\leq\kappa$, if $(\alpha+1,\beta]^{\mathcal{T}}$ does not drop in model then:
      1. i.
        for all $\gamma\in(\alpha,\beta)$, $E^{\mathcal{T}}_{\gamma}$ is the unique $M^{\mathcal{T}}_{\gamma}$-total extender in $E\in\mathbb{E}_{+}(M^{\mathcal{T}}_{\gamma})$ with $\mathrm{cr}(E)=i^{*\mathcal{T}}_{\alpha+1,\gamma}(\mathrm{cr}(E^{\mathcal{T}}_{\alpha}))$, and
      2. ii.
        if $\beta<\kappa$ then:
        1. A.
          $\mathrm{lh}(E^{\mathcal{T}}_{\beta})\leq i^{*\mathcal{T}}_{\alpha+1,\beta}(\mathrm{lh}(E^{\mathcal{T}}_{\alpha}))$ (where if $E^{\mathcal{T}}_{\alpha}=F^{M^{\mathcal{T}}_{\alpha}}$ then this just means that $\mathrm{lh}(E^{\mathcal{T}}_{\beta})\leq\mathrm{lh}(F^{M^{\mathcal{T}}_{\beta}})$, which is true anyway), and
        2. B.
          $\mathrm{lh}(E^{\mathcal{T}}_{\beta})<i^{*\mathcal{T}}_{\alpha+1,\beta}(\mathrm{lh}(E^{\mathcal{T}}_{\alpha}))$ iff $E^{\mathcal{T}}_{\beta}\neq i^{*\mathcal{T}}_{\alpha+1,\beta}(E^{\mathcal{T}}_{\alpha})$ iff $\beta+1\in\mathscr{D}^{\mathcal{T}}$.
Moreover, in case $N\ntrianglelefteq M$, the tree $\mathcal{T}$ mentioned above is uniquely determined.
The following claim is now a straightforward consequence:
Claim 2.
Let $N,N^{\prime}$ be good, with $N,N^{\prime}\ntrianglelefteq M$, and let $\mathcal{T},\mathcal{T}^{\prime}$ be the respective trees as in Claim 1. Then either:
  1. 1.
    $N|\kappa=N^{\prime}|\kappa$ and $\mathcal{T}=\mathcal{T}^{\prime}$ and either $N\trianglelefteq N^{\prime}$ or $N^{\prime}\trianglelefteq N$, or
  2. 2.
    $N|\kappa\neq N^{\prime}|\kappa$, there is some (uniquely determined) $\alpha<\kappa$ such that $\mathcal{T}\!\upharpoonright\!(\alpha+1)=\mathcal{T}^{\prime}\!\upharpoonright\!(\alpha+1)$ and $E^{\mathcal{T}}_{\alpha}\neq E^{\mathcal{T}^{\prime}}_{\alpha}$, and if $\mathrm{lh}(E^{\mathcal{T}}_{\alpha})<\mathrm{lh}(E^{\mathcal{T}^{\prime}}_{\alpha})$, then $\mathcal{T}\!\upharpoonright\![\alpha,\kappa+1)$ can be considered as a tree on $N^{\prime}|\kappa$, with properties like those of the trees in Claim 1, and $\mathcal{T}\!\upharpoonright\![\alpha,\kappa+1)\in N^{\prime}|\gamma^{N^{\prime}}$ and $N\in N^{\prime}|\gamma^{N^{\prime}}$; otherwise $\mathrm{lh}(E^{\mathcal{T}^{\prime}}_{\alpha})<\mathrm{lh}(E^{\mathcal{T}}_{\alpha})$ and it is symmetric.
And the next claim is also a straightforward consequence of the previous two:
Claim 3.
$\in$ wellorders the good premice.
Now given $X\in\mathcal{H}^{M}_{\kappa^{+M}}$, let $N_{X}$ be the $\in$-least good premouse $N$ with $X\in N$. We define a good wellorder $<^{*}$ over $\mathcal{H}^{M}_{\kappa^{+M}}$ by setting $X<^{*}Y$ iff either:
  1. $N_{X}\in N_{Y}$, or
  2. $N_{X}=N_{Y}$ and $X<^{N_{X}}Y$ (where $<^{N_{X}}$ is the order of constructibility of $N_{X}$).
Claim 4.
$<^{*}$ is a $\Sigma_{1}^{\left\lfloor M\right\rfloor}(\{x,\kappa\})$ good wellorder of $\mathcal{H}^{M}_{\kappa^{+M}}$.
Proof.
This is straightforward; just note that if $N,N^{\prime}$ are good and $N^{\prime}\in N$ with $N|\kappa\neq N^{\prime}|\kappa$, then in fact $N^{\prime}\triangleleft M^{\mathcal{T}}_{\kappa}$ for some $\mathcal{T}\in N$ as described above. So $N$ can identify the collection of good premice $N^{\prime}\in N$ by making use of this. Thus, we get that $X<^{*}Y$ iff there are good premice $N,P$ with $X\in N$ and $Y\in P$, and $N$ thinks (using the preceding discussion) that there is no good premouse $N^{\prime}\in N$ with $X\in N^{\prime}$, and likewise $P$ thinks there is no good $P^{\prime}\in P$ with $Y\in P^{\prime}$ (this ensures that $N=N_{X}$ and $P=N_{Y}$), and either $N\in P$, or $N=P$ and $X<^{N}Y$. ∎
Subcase 2.2.
Otherwise.
Fix an $M$-cardinal $\eta$ such that $\delta<\eta<\kappa$, where $\delta$ is the supremum of the $M$-measurables which are $<\kappa$ (by the subcase hypothesis, there are some), and $\eta=\theta^{+M}$ for some $\theta$. Set $x=M|\eta$. Working in $M$, we define good premice as before (from the parameter $M|\eta$), except that we don’t demand that $\eta$ be a cutpoint of $N$.
Claim 5.
Let $N$ be good. Then either:
  1. $N\trianglelefteq M$, or
  2. there is an $\omega$-maximal iteration tree $\mathcal{T}$ on $M$ such that:
    1. (a)
      $\eta<\mathrm{lh}(E^{\mathcal{T}}_{0})$, $\mathcal{T}$ is based on $M|\kappa$, and $\mathrm{lh}(\mathcal{T})=\kappa+1$,
    2. (b)
      $N\trianglelefteq M^{\mathcal{T}}_{\kappa}$,
    3. (c)
      $\mathcal{T}\in M$,
    4. (d)
      if $\mathrm{cr}(E^{\mathcal{T}}_{\alpha})<\eta$ then $\mathrm{pred}^{\mathcal{T}}(\alpha+1)=0$, and if $\mathrm{cr}(E^{\mathcal{T}}_{\alpha})\geq\eta$ then $\mathrm{pred}^{\mathcal{T}}(\alpha+1)=\alpha$,
    5. (e)
      $[0,\alpha+1]^{\mathcal{T}}\cap\mathscr{D}^{\mathcal{T}}=\emptyset$ iff $\mathrm{pred}^{\mathcal{T}}(\alpha+1)=0$ and $\mathrm{cr}(E^{\mathcal{T}}_{\alpha})<\eta$,
    6. (f)
      for every $\alpha+1<\mathrm{lh}(\mathcal{T})$ with $\eta\leq\mathrm{cr}(E^{\mathcal{T}}_{\alpha})$, $M^{\mathcal{T}}_{\alpha}||\mathrm{lh}(E^{\mathcal{T}}_{\alpha})\models$“There are no measurable cardinals $>\delta$”, and
    7. (g)
      for all $\alpha+1\leq\beta\leq\kappa$, if $\eta\leq\mathrm{cr}(E^{\mathcal{T}}_{\alpha})$ and $(\alpha+1,\beta]^{\mathcal{T}}$ does not drop in model then:
      1. i.
        for all $\gamma\in(\alpha,\beta)$, $E^{\mathcal{T}}_{\gamma}$ is the unique $M^{\mathcal{T}}_{\gamma}$-total extender $E\in\mathbb{E}_{+}(M^{\mathcal{T}}_{\gamma})$ with $\mathrm{cr}(E)=i^{*\mathcal{T}}_{\alpha+1,\gamma}(\mathrm{cr}(E^{\mathcal{T}}_{\alpha}))$, and
      2. ii.
        if $\beta<\kappa$ then:
        1. A.
          $\mathrm{lh}(E^{\mathcal{T}}_{\beta})\leq i^{*\mathcal{T}}_{\alpha+1,\beta}(\mathrm{lh}(E^{\mathcal{T}}_{\alpha}))$, and
        2. B.
          if $\mathrm{pred}^{\mathcal{T}}(\beta+1)=\beta$ then
          $\mathrm{lh}(E^{\mathcal{T}}_{\beta})<i^{*\mathcal{T}}_{\alpha+1,\beta}(\mathrm{lh}(E^{\mathcal{T}}_{\alpha}))\iff E^{\mathcal{T}}_{\beta}\neq i^{*\mathcal{T}}_{\alpha+1,\beta}(E^{\mathcal{T}}_{\alpha})\iff\beta+1\in\mathscr{D}^{\mathcal{T}}.$
    8. (h)
      for all $\alpha+1<\kappa$, if $\mu=\mathrm{cr}(E^{\mathcal{T}}_{\alpha})<\eta$ then letting $\mu^{\prime}$ be least such that $\mu^{\prime}\geq\eta$ and $\mu^{\prime}$ is measurable in $M^{\mathcal{T}}_{\alpha+1}$, then $\mu^{\prime}\leq i^{\mathcal{T}}_{0,\alpha+1}(\mu)$, and letting $E\in\mathbb{E}_{+}(M^{\mathcal{T}}_{\alpha+1})$ be the unique extender which is $M^{\mathcal{T}}_{\alpha+1}$-total with $\mathrm{cr}(E)=\mu^{\prime}$, we have:
      1. i.
        $\mathrm{lh}(E^{\mathcal{T}}_{\alpha+1})\leq\mathrm{lh}(E)$, and
      2. ii.
        if $\mathrm{pred}^{\mathcal{T}}(\alpha+2)=\alpha+1$ then
        $\mathrm{lh}(E^{\mathcal{T}}_{\alpha+1})<\mathrm{lh}(E)\iff E^{\mathcal{T}}_{\alpha+1}\neq E\iff\alpha+2\in\mathscr{D}^{\mathcal{T}}.$
Moreover, in case $N\ntrianglelefteq M$, the tree $\mathcal{T}$ mentioned above is uniquely determined.
Proof.
We use an argument from [18] (also see [22]), using a property related to a feature of $K$ (see [14]). Let $(\mathcal{T},\mathcal{U})$ be the comparison of $(M,N)$. The main thing is to see that $\mathcal{U}$ is trivial, and for this, we need to see that $\mathcal{U}$ is above $\eta$. But otherwise, letting $\alpha$ be least (or in fact any ordinal) such that $\mathrm{cr}(E^{\mathcal{U}}_{\alpha})<\eta$, we get that $E^{\mathcal{U}}_{\alpha}\in M$ and $(M^{\mathcal{T}}_{\alpha}||\mathrm{lh}(E^{\mathcal{U}}_{\alpha}),E^{\mathcal{U}}_{\alpha})$ is a premouse. By [22, Theorem 3.8], and since $\mathcal{T}\in M$, it follows that $E^{\mathcal{U}}_{\alpha}\in\mathbb{E}_{+}(M^{\mathcal{T}}_{\alpha})$, so $E^{\mathcal{U}}_{\alpha}$ was not causing a disagreement, a contradiction. ∎
Note also that $\mathcal{T}$ can only use finitely many extenders $E$ with $\mathrm{cr}(E)<\eta$, since otherwise, because of the tree structure described above ($\mathcal{T}$ is linear except for those $E$’s). Thus, we again get that if $N,N^{\prime}$ are good with $N|\kappa\neq N^{\prime}|\kappa$, then letting $\mathcal{T},\mathcal{T}^{\prime}$ be the corresponding trees on $M$, either $\mathcal{T}$ can be converted into an essentially equivalent tree $\mathcal{U}$ on $N^{\prime}$ with $\mathcal{U}\in N^{\prime}$ and $N\triangleleft M^{\mathcal{U}}_{\kappa}$, or vice versa. So the rest is just like in Subcase 2.1.∎
As a corollary of the previous result applied to $M_{1}|\delta^{M_{1}}$, combined with 3.3:
5.3 Corollary.
Suppose $M_{1}$ exists and is fully iterable. Then there is a proper class inner model satisfying ZFC + “there is a Woodin cardinal” + “for every uncountable cardinal $\kappa$:
  1. there is a good wellorder of $\mathcal{H}_{\kappa^{+}}$, and
  2. the following are equivalent:
    1. (a)
      $\kappa$ is not a limit of measurable cardinals,
    2. (b)
      there is a $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ wellorder of a subset of $\mathcal{P}(\kappa)$ of cardinality $>\kappa$,
    3. (c)
      there is a $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ wellorder of $\mathcal{H}_{\kappa^{+}}$,
    4. (d)
      there is a $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ good wellorder of $\mathcal{H}_{\kappa^{+}}$”.
In [10], it is shown that if $M_{1}^{\#}(x)$ exists for all $x\subseteq\omega_{1}$, then there is no $\Sigma_{1}(\{\omega_{1}\})$ wellorder of $\mathcal{P}(\omega)$. In every proper class $1$-small mouse $M$, the standard $M$ order of construction, restricted to $\mathbb{R}^{M}$, is $\Sigma_{1}(\{\omega_{1}\})$. So by the following fact, the least proper class mouse $M$ where this fails is $L[M_{1}^{\#}]$:
5.4 Theorem.
In $L[M_{1}^{\#}]$, there is no $\Sigma_{1}(\{\omega_{1}\})$ wellorder of $\mathcal{P}(\omega)$.
Proof.
It is well known that $\omega_{1}^{L[M_{1}^{\#}]}$ is measurable in $\mathrm{HOD}^{L[M_{1}^{\#}]}$ (see [16, Remark 3.2***] for a proof). Note that this implies the theorem immediately. ∎
5.5 Remark.
In [18], [22] and [19], there are techniques described via which, in various settings, a mouse $M$ can identify its own internal extender sequence $\mathbb{E}^{M}$ definably over its universe $\left\lfloor M\right\rfloor$. Using some of those results, we get that in short extender mice $M$, if $\kappa<\mathrm{OR}^{M}$ is an $M$-cardinal and either $M\models$$\kappa^{+}$ exists” or $M\models$ ${}^{-}$, the usual $M$-order of constructibility is $\Delta_{4}^{\mathcal{H}^{M}_{\kappa^{+M}}}(\{x\})$ where $x=M|\omega_{1}^{M}$ (that is, $x$ is the initial segment of $M$ through to $\omega_{1}^{M}$, including its extender sequence there (which consists of partial extenders only). Moreover, in many cases, we can make do with $\Delta_{2}$ instead of $\Delta_{4}$. The author does not know whether one can in fact always make do with $\Delta_{2}$. (See [19].) If $M$ is tame, the parameter $M|\omega_{1}^{M}$ can be replaced with a real (see [16]).
5.6 Remark.
Jouko Väänänen and Philip Welch argue in [27] that if $M$ is a proper class 1-small mouse satisfying “there is no proper class model with a Woodin cardinal” then in $\left\lfloor M\right\rfloor$, the relation “$x=\mathcal{P}(y)$” is $\Sigma_{1}(\mathrm{Card})$. Their argument uses facts about the core model $K$ below 1 Woodin cardinal, and that $M=K^{M}$. It uses in particular [27, Lemma 3.9] (which as literally written in [4] also assumes a large enough measurable cardinal $\Omega$, which might not be available here; in [27] it is pointed out that this is not needed, given the more recent development of the theory of $K$ at this level without the measurable cardinal). Using instead an argument like that in the proof of Claim 5 of the proof of Theorem 5.1, one can prove the Welch-Väänänen result just mentioned without referring to $K$. That is, work in a proper class 1-small mouse $M$ which satisfies “there is no proper class inner model with a Woodin cardinal” and let $N$ be a mouse such that $\mathrm{OR}^{N}=\lambda^{++}$ for some $\lambda$, and $\mathrm{Card}^{N}=\mathrm{Card}\cap N$. (As discussed there, the iterability of $N$ can be expressed in an appropriate manner.) We want to see that $N\trianglelefteq M$. So let $(\mathcal{T},\mathcal{U})$ be the comparison of $(M|\lambda^{++},N)$, with both trees at degree 00. We want to see that $\mathcal{T},\mathcal{U}$ are both trivial. Suppose not. Then easily, $\mathcal{T}$ cannot be trivial. Let $\gamma=\mathrm{lh}(E^{\mathcal{T}}_{0})$ or $\gamma=\mathrm{lh}(E^{\mathcal{U}}_{0})$, for whichever of these extenders is non-empty, and let $\eta=\mathrm{card}^{M}(\gamma)$. Then $\mathcal{U}$ is above $\eta$, like in the proof of Claim 5. But then $\mathcal{U}$ drops along $b^{\mathcal{U}}$ (since $\gamma<\eta^{+M}=\eta^{+N}$). So $b^{\mathcal{T}}$ does not drop, and $M^{\mathcal{T}}_{\infty}\triangleleft M^{\mathcal{U}}_{\infty}$. Standard weasel arguments now show that $\mathrm{lh}(\mathcal{T},\mathcal{U})=\lambda^{++}+1$ and $\mathrm{OR}(M^{\mathcal{T}}_{\lambda^{++}})=\lambda^{++}<\mathrm{OR}(M^{\mathcal{U}}_{\lambda^{++}})$, and for club many such $\beta<\lambda^{++}$, we have $\beta<^{\mathcal{U}}\lambda^{++}$ and $\beta=\mathrm{cr}(i^{\mathcal{U}}_{\beta\lambda^{++}})$ and $i^{\mathcal{U}}_{\beta\lambda^{++}}(\beta)=\lambda^{++}$. But then $M^{\mathcal{T}}_{\lambda^{++}}$ has largest cardinal $i^{\mathcal{T}}_{0\lambda^{++}}(\lambda^{+})$, whereas in $M^{\mathcal{U}}_{\lambda^{++}}$, $\lambda^{++}$ is inaccessible, a contradiction.
In case $\kappa$ is a limit cardinal of $M$, the good wellorder of $(\mathcal{H}_{\kappa^{+}})^{M}$ defined in the proof of Theorem 5.1 is not in general the same as the usual order of constructibility. In fact, if $M|\kappa$ is closed under sharps, then they disagree, because the set $M|\kappa$ itself is positioned strictly above sets of higher $V$-rank which appear in the dropping iterates of $M|\kappa$ which arise in the proof. So one can ask whether the usual order of constructibility could be $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ in $M$. We show below that if $M=M_{1}$ and $\kappa$ is $\omega_{1}$-iterable in $M_{1}$, then it is not, and in fact:
5.7 Theorem.
Work in $M_{1}$. Let $\kappa$ be an uncountable cardinal of $M_{1}$ which is not a limit of measurables in $M_{1}$. Then:
  1. 1.
    If $M_{1}\models$$\kappa$ is not Mahlo” then $M_{1}|\kappa$ is $\Delta_{1}(\mathcal{H}_{\kappa}^{M_{1}}\cup\{\kappa\})$.
  2. 2.
    If $M_{1}\models$$\kappa$ is $\omega_{1}$-iterable” then $M_{1}|\kappa$ is not $\Delta_{1}(\mathcal{H}_{\kappa}^{M_{1}}\cup\{\kappa\})$.
Proof.
Part 1: If $\kappa$ is a successor cardinal use the proof of Theorem 5.1. So suppose $\kappa$ is a limit cardinal. Fix an $M_{1}$-cardinal $\eta<\kappa$ as in the proof of Theorem 5.1.
Suppose first that $\kappa$ is singular. Then $M_{1}|\kappa$ is the unique premouse $P$ such that $M_{1}|\eta\triangleleft P$, $\mathrm{OR}^{P}=\kappa$, $P\models$“there are no measurable cardinals $\geq\eta$”, there is a sound premouse $Q$ such that $P\triangleleft Q$, $\rho_{\omega}^{Q}=\kappa$, $Q\models$$\kappa$ is singular”, and there is $\xi$ such that $\mathcal{J}_{\xi}[Q]\models$$Q$ is strongly iterable” (compare $M_{1}$ with $Q$, and note that because $\kappa$ is singular in $Q$, we can’t have that the $M_{1}$ side drops on its main branch).
Now suppose $\kappa$ is inaccessible, but not Mahlo. So there is a club $C\subseteq\kappa$ such that $\beta$ is singular for all $\beta\in C$. So there is a sequence $\vec{f}=\left<f_{\beta}\right>_{\beta\in C}$ of witnessing singularizing functions. We can have our $\Delta_{1}$ description be much as before, but instead $Q\models$$\kappa$ is singular”, $Q\models$$\kappa$ is not Mahlo, as witnessed as by some $(C,\vec{f})$”. Now argue much like before.
Part 2: Suppose $\kappa$ is $\omega_{1}$-iterable but $M_{1}|\kappa$ is $\Delta_{1}(\mathcal{H}_{\kappa}^{M_{1}}\cup\{\kappa\})$. So $\{M_{1}|\kappa\}$ is $\Sigma_{1}(\mathcal{H}_{\kappa}^{M_{1}}\cup\{\kappa\})$. Let $\varphi$ be $\Sigma_{1}$ and $p\in M_{1}|\kappa$ be such that $M_{1}|\kappa$ is the unique $X\in M_{1}$ such that $M_{1}\models\varphi(p,\kappa,X)$. So there is some $\alpha<\kappa^{+M_{1}}$ such that $M_{1}|\alpha\models\varphi(p,\kappa,X)$. Let $\beta\in(\alpha,\kappa^{+M_{1}})$ be such that $M_{1}|\beta\models$ ${}^{-}+$$\kappa$ is the largest cardinal” and $U$ be some $\omega_{1}$-iterable weakly amenable ultrafilter over $M_{1}|\beta$. Let
$\pi:(\bar{M},\bar{U})\to(M_{1}|\beta,U)$
be elementary with $\bar{\kappa}=\mathrm{card}(\bar{M})<\kappa$ and $\pi(\bar{\kappa})=\kappa$ and $p\in M_{1}|\bar{\kappa}$. Then $(\bar{M},\bar{U})\models\varphi(p,\bar{\kappa},M_{1}|\bar{\kappa})$, $\bar{U}$ is weakly amenable to $\bar{M}$ and $(\bar{M},\bar{U})$ is $\omega_{1}$-iterable. Let $(\bar{M}^{\prime},\bar{U}^{\prime})$ be the $\kappa$th iterate and $i:(\bar{M},\bar{U})\to(\bar{M}^{\prime},\bar{U}^{\prime})$ the iteration map. Then $i$ is elementary as a map $\bar{M}\to\bar{M}^{\prime}$ and $i(\bar{\kappa})=\kappa$, so $\bar{M}^{\prime}\models\varphi(p,\kappa,\bar{M}^{\prime}|\kappa)$, so in fact $\varphi(p,\kappa,\bar{M}^{\prime}|\kappa)$ holds, but $\bar{M}^{\prime}|\kappa\neq M_{1}|\kappa$, a contradiction. ∎
6 Weak compactness and the perfect set property
We now answer [9, Question 10.1]. The answer is “yes”:
6.1 Theorem.
Assume ZFC and let $\kappa$ be a weakly compact cardinal such that for every $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\{\kappa\})$ set $D\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$, there is a perfect embedding $\iota:{{}^{\kappa}}\kappa\to\mathcal{P}(\kappa)$ with $\mathrm{rg}(\iota)\subseteq D$. Then there is a proper class inner model satisfying “there is a weakly compact limit of measurable cardinals”.
Proof.
[…]
Claim 1.
$\kappa$ is weakly compact in $K$.
Proof.
Let $T\in K$ be a tree on $\kappa$ of the relevant form. Using the weak compactness of $\kappa$ in $V$, let $j:M\to N$ be elementary, where $M,N$ are transitive, $V_{\kappa}\cup\{\kappa,T\}\in M$, for some $\alpha<\kappa^{+K}$, we have $T\in K|\alpha\in M$, $M\models\mathsf{ZFC}^{-}$+“$K|\alpha$ is iterable” and ${{}^{<\kappa}}M\subseteq M$, with $\mathrm{cr}(j)=\kappa$. We may assume that $N=\mathrm{Ult}_{0}(M,U)$ where $U$ is the normal measure derived from $j$, and therefore ${{}^{<\kappa}}N\subseteq N$. Let $b\in j(T)$ be a branch at level $\kappa$ of $j(T)$. Then $b\subseteq T$, so it suffices to see that $b\in K$. But $b\subseteq\kappa$ and $b\in j(K|\kappa)$, so $b\in j(K|\kappa)|\beta$ for some $\beta<\kappa^{+j(K|\kappa)}$ such that $j(K|\kappa)|\beta$ projects to $\kappa$. And $N\models$$j(K|\kappa)$ is iterable”, so $j(K|\kappa)$ really is iterable, since ${{}^{<\kappa}}N\subseteq N$. But then considering that $K$ is below a measurable limit of measurables, it follows that $j(K|\kappa)|\beta\triangleleft K$, which suffices. ∎
By the claim, it suffices to see that $\kappa$ is limit of measurables in $K$. So suppose we can fix an inaccessible cardinal $\eta<\kappa$ such that all measurables of $K$ below $\kappa$ are $<\eta$.
Claim 2.
$\kappa^{+K}=\kappa^{+}$.
Proof.
Suppose not. Then by the weak compactness of $\kappa$, we can fix some transitive set $M$ of cardinality $\kappa$ with $\mathcal{P}(\kappa)\cap K\subseteq M$ and ${{}^{\omega}}M\subseteq M$, another transitive $N$, and an elementary $j:M\to N$ with $\mathrm{cr}(j)=\kappa$, and such that if $\kappa$ is measurable in $K$, then the (unique) normal measure $D$ on $\kappa$ in $K$ is also in $M$, and in fact $D$ is coded by a subset of $\kappa$ in $M$, and hence $D\in N$ also in this case; we may also assume that $M\models$ ${}^{-}$ and that $N=\mathrm{Ult}_{0}(M,U)$ where $U$ is the normal $M$-ultrafilter over $\kappa$ derived from $j$, so then ${{}^{\omega}}N\subseteq N$.
Let $U$ be the normal measure over $K$ derived from $j$. Let $K^{\prime}=\mathrm{Ult}(K,U)$. Then since ${{}^{\omega}}M\subseteq M$, $K^{\prime}$ is wellfounded. (If $f_{n}\right>_{n<\omega}\subseteq K$ with $f_{n}:\kappa\to\mathrm{OR}$ and $A_{n}=\{\alpha<\kappa\bigm{|}f_{n+1}(\alpha)<f_{n}(\alpha)\}\in U$ for all $n<\omega$, then since $\vec{A}=\left<A_{n}\right>_{n<\omega}\in M$, we get $j(\vec{A})\in N$, but $\mu\in j(A_{n})$ for each $n$, so $\bigcap j(\vec{A})\neq\emptyset$, so $\bigcap\vec{A}\neq\emptyset$, which gives a contradiction.)
So we have an elementary $\ell:K\to K^{\prime}$ with $\mathrm{cr}(\ell)=\kappa$. So by [23, §8], $K^{\prime}$ is a normal iterate of $K$ and $\ell$ is the iteration map. Note then that $U=D\in K$ is the unique normal measure on $\kappa$ in $K$, and $\ell=i^{K}_{D}$ is the ultrapower map. So $U=D\in M\cap N$. But $\mathcal{P}(\kappa)\cap K=\mathcal{P}(\kappa)\cap j(K)$, like in the proof of Claim 1. We have $D\in N$, and note that $\mathrm{Ult}(j(K),D)$ is also wellfounded. We may also assume that $M,N$ themselves satisfy enough of [23, §8] that it follows that $D\in j(K)$, so $j(\kappa)$ is measurable in $j(K)$, so $\kappa$ is a limit of measurables of $K$, contradiction. ∎
Recall we fixed $\eta$ earlier, bounding the measurables of $K$ which are $<\kappa$. We may assume $\eta=\theta^{+K}$ for some $\theta$. Now let $D$ be the set of subsets of $\kappa$ coding a pair $(M,T)$, where $M$ is an iterable premouse of height $\kappa$ such that $K|\eta\triangleleft M$, $M$ has no measurable cardinals $\geq\eta$, there is a sound premouse $P$ such that $M\triangleleft P$, $\rho_{\omega}^{P}=\kappa$, $P$ satisfies the condensation theorem (see [25]) and $T$ is the theory of $P$ in parameters in $\kappa$. Note that $D$ is $\Sigma_{1}(\{K|\eta,\kappa\})$, and $\mathrm{card}(D)>\kappa$, since $\kappa^{+K}=\kappa^{+}$ and for cofinally many proper segments $P\triangleleft K|\kappa^{+K}$, we have $(K|\kappa,T_{P})\in D$ where $T_{P}=\operatorname{Th}^{P}(\kappa)$. So we can fix a perfect function $\iota:{{}^{\kappa}}\kappa\to\mathcal{P}(\kappa)$ with $\mathrm{rg}(\iota)\subseteq D$.
Claim 3.
There are at most $\kappa$-many elements of $D$ of form $(M,T)$ for some $M\neq K|\kappa$.
Proof.
Fix such a pair $(M,T)$. The comparison of $K|\kappa$ versus $M$, producing iterations $\mathcal{T}$ on $K|\kappa$ and $\mathcal{U}$ on $M$, is such that $b^{\mathcal{U}}$ does not drop, but $b^{\mathcal{T}}$ drops and $M^{\mathcal{U}}_{\infty}\triangleleft M^{\mathcal{T}}_{\infty}$. (Otherwise we would have that neither $b^{\mathcal{T}}$ nor $b^{\mathcal{U}}$ drops, and $M^{\mathcal{T}}_{\infty}=M^{\mathcal{U}}_{\infty}$, but then both sides are trivial, because $K$ has no measurables $\geq\eta$, and so $M=K|\kappa$, contradiction.) So in fact, $\mathcal{U}$ is trivial, since $M$ has no measurables $\geq\eta$, and $K|\eta=M|\eta$. It follows that after some stage $\alpha<\kappa$, $\mathcal{T}$ simply consists of linearly iterating some fixed measure out to $\kappa$, producing $M=M^{\mathcal{T}}_{\kappa}|\kappa$. So there are at most $\kappa$-many such $M$. ∎
So by diagonalizing if necessary, we may assume that all elements of $\mathrm{rg}(\iota)$ are of form $(K|\kappa,T)$. Note that also, letting $(M,T)\in D$ with $M=K|\kappa$ and $P$ the model with theory $T$, we have $P\triangleleft K$, by Jensen’s condensation argument. Now consider $\iota$ as a forcing $\mathbb{P}_{\iota}$, and let $G$ be $(V,\mathbb{P}_{\iota})$-generic. Like in the proof of Theorem 4.1 part 5, this produces a contradiction. ∎
7 Almost disjoint families
We summarized our results on almost disjoint families in §1.
7.1 Theorem.
Assume ZFC and let $\kappa$ be an uncountable cardinal. Then:
  1. 1.
    If $\operatorname{cof}(\kappa)=\omega$ then there is a $\Delta_{1}(\{\kappa\})$ almost disjoint family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$, such that for all $A\in\mathscr{F}$, $A$ is unbounded in $\kappa$ and has ordertype $\omega$.
  2. 2.
    Suppose $\kappa$ is a limit of measurable cardinals. Then:
    1. (a)
      There is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\geq\kappa$.
    2. (b)
      If $\operatorname{cof}(\kappa)>\omega$ then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ almost disjoint family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$.
  3. 3.
    Let $\mu<\kappa$ and suppose that $\mu$ is measurable and $\kappa$ is $\mu$-steady. Let $U$ be a normal measure on $\mu$. Let $S=\{\alpha\in\mathrm{OR}\bigm{|}i^{\mathcal{T}_{U}}_{0\lambda}(\alpha)=\alpha\text{ for all }\lambda<\kappa\}$ (so $\kappa,\kappa^{+}\in S$). Then there is no $\Sigma_{1}(\mathcal{H}_{\mu}\cup S)$ mad family $\mathscr{F}$ at $\kappa$ of cardinality $>2^{\mu}$.
Proof.
Part 1 is routine: just let $A\in\mathscr{F}$ iff $A\subseteq\kappa$, $A$ has ordertype $\omega$, $A$ is unbounded in $\kappa$ and for every $\alpha\in A$, $\alpha$ is the Gödel code of a pair $(\beta,\gamma)$ such that $\beta$ is the Gödel code of $A\cap\alpha$. This easily works.
Part 2b: Suppose otherwise and fix $\vec{p}\in\mathcal{H}_{\kappa}\cup\mathrm{OR}$ and some $\Sigma_{1}(\{\vec{p}\})$ almost disjoint family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $>\kappa$. Let $\eta=\operatorname{cof}(\kappa)$ and let $\kappa_{\alpha}\right>_{\alpha<\eta}$ be a strictly increasing sequence of measurable cardinals cofinal in $\kappa$, with $\eta<\kappa_{0}$ if $\eta<\kappa$, and $\vec{p}$ fixed by all iterations of measures which are based on $V_{\delta}$ for some $\delta<\kappa$, are above $\kappa_{0}$, and have length $<\kappa$. (Such a $\kappa_{0}$ exists by Lemma 2.11.) Let $U_{\alpha}\right>_{\alpha<\eta}$ a sequence with $U_{\alpha}$ a normal measure on $\kappa_{\alpha}$. Since $\mathscr{F}$ has cardinality $>\kappa$, we can fix $X\in\mathscr{F}$ such that $i^{\mathcal{T}_{U_{\alpha}}}_{0\kappa}(X\cap\kappa_{\alpha})\neq X$ for all $\alpha<\eta$. By thinning out the sequence, we may assume that $i^{\mathcal{T}_{U_{\alpha}}}_{0\kappa_{\alpha+1}}(X\cap K_{\alpha})\neq X\cap\kappa_{\alpha+1}$ for every $\alpha<\eta$. Let $\lambda=\sup_{n<\omega}\kappa_{n}$; so $\lambda<\kappa$. For $\xi\in[\lambda,\kappa)$ let $n_{\xi}$ be the least $n<\omega$ such that $i^{\mathcal{T}_{U_{n}}}_{0\kappa_{n+1}}(\xi)=\xi$; again by Lemma 2.11, such an $n$ exists. Then since $\operatorname{cof}(\kappa)>\omega$ and $X$ is unbounded in $\kappa$, there is $n<\omega$ and an unbounded set $Y\subseteq X$ such that $n_{\xi}=n$ for all $\xi\in Y$. Fix such $n,Y$. Let $j=i^{\mathcal{T}_{U_{n}}}_{0\kappa_{n+1}}$. Then $j(X)\in\mathscr{F}$ since $j$ fixes $\vec{p}$. And $j(X)\neq X$ by construction. But $Y\subseteq X\cap j(X)$, and since $Y$ is unbounded in $\kappa$, this is a contradiction.
Part 3: Suppose otherwise. Fix $\vec{p}\in\mathcal{H}_{\mu}\cup S$ and a $\Sigma_{1}(\{\vec{p}\})$ mad family $\mathscr{F}$ at $\kappa$ of cardinality $>2^{\mu}$.
We have $i^{\mathcal{T}_{U}}_{0\lambda}(\mathscr{F})=\mathscr{F}\cap M^{\mathcal{T}_{U}}_{\lambda}$ for all $\lambda<\kappa$. For $i^{\mathcal{T}_{U}}_{0\lambda}(\mathscr{F})\subseteq\mathscr{F}$ since $i^{\mathcal{T}_{U}}_{0\lambda}(\vec{p})=\vec{p}$ and by $\Sigma_{1}$ upward absoluteness, and conversely, if $X\in\mathscr{F}\cap M^{\mathcal{T}_{U}}_{\lambda}$, then by the maximality of $\mathscr{F}$ and elementarity, there is $Y\in i^{\mathcal{T}_{U}}_{0\lambda}(\mathscr{F})$ such that $Y\cap X$ is unbounded in $\kappa$, but as just mentioned, it follows that $Y\in\mathscr{F}$ also, and therefore $Y=X$.
For each $\xi<\kappa$ let $\mu_{\xi}=i^{\mathcal{T}_{U}}_{0\xi}(\mu)$. For $\alpha<\kappa$ let $\gamma_{\alpha}$ be the least $\gamma$ such that $\mu_{\gamma}>\alpha$, and choose some $n_{\alpha}<\omega$ and $f_{\alpha}:[\mu]^{n_{\alpha}}\to\mu$ such that for some $c=\{c_{0}<\ldots<c_{n_{\alpha}-1}\}\in[\gamma_{\alpha}]^{<\omega}$, we have $\alpha=i^{\mathcal{T}_{U}}_{0\gamma_{\alpha}}(f_{\alpha})(b_{\alpha})$, where $b_{\alpha}=\{\mu_{c_{0}},\ldots,\mu_{c_{n_{\alpha}-1}}\}$. Note then that $\alpha=i^{\mathcal{T}_{U}}_{0\gamma}(f_{\alpha})(b_{\alpha})$ for all $\gamma\in[\gamma_{\alpha},\kappa]$.
Let us say that $\mu^{\prime}<\kappa$ is nicely stable iff it satisfies the properties of Lemma 2.2 part 2; that is, $i^{\mathcal{T}_{U}}_{0\mu^{\prime}}(\mu)=\mu^{\prime}$ and $i^{\mathcal{T}_{U}}_{0\lambda}(\mu^{\prime})=\mu^{\prime}$ for all $\lambda<\mu^{\prime}$. So the nicely stable ordinals are unbounded in $\kappa$.
Now fix $X\in\mathscr{F}$ such that $X\neq i^{\mathcal{T}_{U}}_{0\kappa}(X\cap\mu)$; such an $X$ exists because $\mathscr{F}$ has cardinality $>2^{\mu}$. Let $\eta=\operatorname{cof}(\kappa)$ and fix a normal function $g:\eta\to\kappa$ with $\mu\leq g(0)$ (but $\kappa$ might be regular). Define a sequence $\alpha_{\beta},A_{\beta}\right>_{\beta<\eta}$ as follows: Let $\alpha_{0}$ be the least $\alpha\in X$ such that $\alpha>g(0)$. Let $A_{0}$ be the set of the least $(n_{\alpha_{0}}+1)$-many nicely stable ordinals which are $>\gamma_{\alpha_{0}}$. Given $\alpha_{\beta},A_{\beta}\right>_{\beta<\lambda}$ where $\lambda<\eta$, let $\alpha_{\lambda}$ be the least $\alpha\in X$ such that $\alpha>g(\lambda)$ and $\alpha>\bigcup_{\beta<\lambda}A_{\beta}$, and let $A_{\lambda}$ be the set of the least $(n_{\alpha_{\lambda}}+1)$-many nicely stable ordinals which are $>\gamma_{\alpha_{\lambda}}$.
Now define, for each $\beta<\eta$,
$\alpha^{\prime}_{\beta}=i^{\mathcal{T}_{U}}_{0\max(A_{\beta})}(f_{\alpha_{\beta}})(A_{\beta}\backslash\{\max(A_{\beta})\}),$
so also
$\alpha^{\prime}_{\beta}=i^{\mathcal{T}_{U}}_{0\kappa}(f_{\alpha_{\beta}})(A_{\beta}\backslash\{\max(A_{\beta})\}).$
Then $\alpha_{\beta}\leq\alpha^{\prime}_{\beta}<\max(A_{\beta})$ for all $\beta<\eta$. Let $Y=\{\alpha^{\prime}_{\beta}\bigm{|}\beta<\eta\}$. So $Y$ is cofinal in $\kappa$, so by the maximality of $\mathscr{F}$, we can fix $Z\in\mathscr{F}$ with $Y\cap Z$ cofinal in $\kappa$.
We claim that $Z\neq i^{\mathcal{T}_{U}}_{0\kappa}(Z\cap\mu)$. For suppose otherwise. Since $X\neq i^{\mathcal{T}_{U}}_{0\kappa}(X\cap\mu)$ but $X,Z\in\mathscr{F}$, $X\cap Z$ is bounded in $\kappa$. So we can fix $\beta$ such that $\alpha_{\beta}\notin Z$ but $\alpha_{\beta}^{\prime}\in Z$. (That is, for all sufficiently large $\beta$, we have $\alpha_{\beta}\notin Z$, since $\alpha_{\beta}\in X$. But $Y\cap Z$ is cofinal in $\kappa$, so there is such a $\beta$ with $\alpha_{\beta}^{\prime}\in Z$.) But because $Z=i^{\mathcal{T}_{U}}_{0\kappa}(Z\cap\mu)$ and $\alpha_{\beta}=i^{\mathcal{T}_{U}}_{0\kappa}(f_{\alpha_{\beta}})(b_{\alpha_{\beta}})$ and $\alpha_{\beta}^{\prime}=i^{\mathcal{T}_{U}}_{0\kappa}(f_{\alpha_{\beta}})(A_{\beta}\backslash\{\max(A_{\beta})\})$, in fact $\alpha_{\beta}\in Z$ iff $\alpha_{\beta}^{\prime}\in Z$, a contradiction.
So there is $\lambda<\kappa$ such that $W=i^{\mathcal{T}_{U}}_{0\lambda}(Z)\neq Z$, but $W\in\mathscr{F}$. So $W\cap Z$ is bounded in $\kappa$. So we can find $\beta<\eta$ such that $\alpha_{\beta}^{\prime}\in Z\backslash W$ and $\lambda<\min(A_{\beta})$. But every ordinal in $A_{\beta}$ is nicely stable, so $i^{\mathcal{T}_{U}}_{0\lambda}(A_{\beta})=A_{\beta}$, from which it easily follows that $i^{\mathcal{T}_{U}}_{0\lambda}(\alpha_{\beta}^{\prime})=\alpha_{\beta}^{\prime}$, so $\alpha_{\beta}^{\prime}\in W$, a contradiction.
Part 2a: This is an easy corollary of part 3 and Lemma 2.11. ∎
Let $M$ be a premouse. If $M\models$$\omega_{1}$ exists” then $\mathrm{HC}^{M}$ is the universe of $M|\omega_{1}^{M}$, but $M|\omega_{1}^{M}$ itself (which has $\mathbb{E}^{M}\!\upharpoonright\!\omega_{1}^{M}$ as a predicate) need not be definable from parameters over $\mathrm{HC}^{M}$. In the following theorem, we might have $\kappa=\omega_{1}^{M}$, in which case $\mathfrak{m}=M|\omega_{1}^{M}\in\mathcal{H}^{M}_{\omega_{2}^{M}}\backslash\mathcal{H}^{M}_{\omega_{1}^{M}}$. But if $\omega_{1}^{M}<\kappa$ then $\mathfrak{m}\in\mathcal{H}_{\kappa}^{M}$.
Recall that a premouse is non-tame if there is $E\in\mathbb{E}^{M}_{+}$ and some $\delta\in[\mathrm{cr}(E),\nu(E)]$ such that $M|\mathrm{lh}(E)\models$$\delta$ is Woodin”; otherwise it is tame. Tame mice can satisfy “there is a strong cardinal which is a limit of Woodin cardinals” and more. They cannot satisfy “there is a Woodin limit of Woodin cardinals”.
7.2 Theorem.
Let $M$ be a $(0,\omega_{1}+1)$-iterable premouse, and let $\kappa\in\mathrm{OR}^{M}$ be such that $M\models$$\kappa>\omega$ is a regular cardinal and $\kappa^{+}$ exists”. Let $\mathfrak{m}=M|\omega_{1}^{M}$. Then:
  1. 1.
    $M\models$“there is a mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa,\mathfrak{m}\})$”.
  2. 2.
    If $M$ is tame then $M\models$“there is a mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa,x\})$ for some $x\in\mathbb{R}$”.
  3. 3.
    If $M=M_{n}$ for some $n<\omega$ (the minimal proper class mouse with $n$ Woodin cardinals) or $M$ is the minimal proper class mouse satisfying “there is a Woodin limit of Woodin cardinals” then $M\models$“there is a mad familiy $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa\})$”.
7.3 Remark.
There are many other examples of “$\varphi$-minimal” mice for which the same conclusion as in part 3 holds.
Proof.
Part 1: By [19], $M|\kappa$ is definable over $\mathcal{H}_{\kappa}^{M}$ from the parameter $\mathfrak{m}$ (or predicate $\mathfrak{m}$, in case $\kappa=\omega_{1}^{M}$, but in this case it is trivial). And working in $M$, where $\kappa$ is regular, $M|\kappa^{+}$ is simply definable from the parameter $M|\kappa$, as the Jensen stack over $M|\kappa$ [^9] That is, it is the stack of all sound premice $P$ such that $M|\kappa\triangleleft P$ and $\rho_{\omega}^{P}=\kappa$ and condensation holds for $P$; we don’t need to assert any iterability about $P$ here.
Now work in $M$. We will define a mad family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ as desired. We will use the standard $M$-constructibility order to guide a recursive construction to build $\mathscr{F}$. Let us sketch how we will ensure that this will be $\Pi_{1}(\{\kappa,\mathfrak{m}\})$ definable. Given a set $X$ of ordinals, and an ordinal $\alpha$, let $\alpha\oplus X=\{\alpha+\beta\bigm{|}\beta\in X\}$. For sets $X,A\subseteq\mathrm{OR}$ and $\alpha\in\mathrm{OR}$, say that $X$ is directly encoded into $A$ at $\alpha$ if there is $\beta\in\mathrm{OR}$ such that $\alpha+X=[\alpha,\beta)\cap A$; say that $X$ is directly encoded into $A$ if there is some such $\alpha$. The key to the construction will be to ensure that for each $A\in\mathscr{F}$, every bounded set $X\subseteq\kappa$ is directly encoded into $A$. Then from any $A\in\mathscr{F}$, we can easily recover $\mathcal{H}_{\kappa}^{M}$, and so by the previous paragraph, recover $M|\kappa$ and (proper segments of) $M|\kappa^{+}$, and hence the recursive construction used to build $\mathscr{F}$.
So fix a surjection $\pi:\kappa\to\mathcal{P}({<\kappa})$ which is definable over $\mathcal{H}_{\kappa}^{M}$ from the parameter $\mathfrak{m}$ and such that $\pi(\alpha)\subseteq\alpha$ for each $\alpha<\kappa$. We will define a sequence $A_{\alpha}\right>_{\alpha<\kappa^{+}}$ by recursion on $\alpha$, and set =$\mathscr{F}=\{A_{\alpha}\}_{\alpha<\kappa^{+}}$. We will also define a sequence $C_{\alpha}\right>_{\alpha<\kappa^{+}}$ of clubs $C_{\alpha}\subseteq\kappa$.
Let us start with $A_{0}$ and $C_{0}$. We define $A_{0}$ through a $\kappa$-sequence of stages $\alpha<\kappa$, at stage $\alpha$ extending with a certain empty interval, followed by a direct encoding of $\pi(\alpha)$. Recall that for a set $X$ of ordinals, the strict supremum $\mathrm{strsup}(X)$ of $X$ is the least ordinal $\eta$ such that $X\subseteq\eta$. Define a sequence $B_{\alpha}\right>_{\alpha<\kappa}$ of sets as follows. Set $B_{0}=\emptyset$. Given $B_{\alpha}$, let $B_{\alpha+1}=B_{\alpha}\cup((\beta+\beta)\oplus\pi(\alpha))$ where $\beta=\mathrm{strsup}(B_{\alpha})$. (We only really need “$\beta+\beta$” instead of just “$\beta$” in case $\alpha$ is a limit ordinal with $\alpha=\mathrm{strsup}(B_{\alpha})=\sup(B_{\alpha})$.) Given $B_{\alpha}$ for $\alpha<\lambda$, where $\lambda$ is a limit, let $B_{\lambda}=\bigcup_{\alpha<\lambda}B_{\alpha}$. Now set $A_{0}=\bigcup_{\alpha<\kappa}B_{\alpha}$. Also let $C_{0}\subseteq\kappa$ be the club of all limit ordinals $\alpha<\kappa$ such that $\alpha=\sup_{\beta<\alpha}B_{\beta}$.
Suppose we have defined $A_{\alpha},C_{\alpha}\right>_{\alpha<\lambda}$ where $0<\lambda<\kappa^{+M}$, and the following conditions hold, for all $\alpha<\kappa$:
  1. (i)
    $A_{\alpha}$ is unbounded in $\kappa$,
  2. (ii)
    every bounded subset of $\kappa$ is directly encoded into $A_{\alpha}$,
  3. (iii)
    $A_{\beta}\cap A_{\alpha}$ is bounded in $\kappa$ for all $\beta<\alpha$,
  4. (iv)
    $C_{\alpha}\subseteq\kappa$ is club and $[\eta,\eta+\eta)\cap A_{\alpha}=\emptyset$ for all $\eta\in C_{\alpha}$.
Now let $B\subseteq\kappa$ be $<^{M}$-least such that $B$ is unbounded in $\kappa$ but there is no $\alpha<\lambda$ such that $B\cap A_{\alpha}$ is unbounded in $\kappa$, if there is such a $B$; otherwise we are finished. We will define $A_{\lambda}$, ensuring that $B\cap A_{\lambda}$ is unbounded in $\kappa$, and maintaining the conditions (i)(iv), and ensure also a bit more.
Fix the $<^{M}$-least surjection $\sigma:\kappa\to\lambda$. We construct $A_{\lambda}$ in $\kappa$-many stages $\alpha<\kappa$. At successor stages $\alpha+1$, we find a suitable point at which to concatenate the set $\pi(\alpha)$, followed by some point in $B$, to what we have produced so far. At limit stages, we insert some space, followed by some point in $B$. So, we define an increasing sequence of sets $a_{\alpha}\right>_{\alpha<\kappa}$; each $a_{\alpha}$ will be an initial segment of $A_{\lambda}$. Let $a_{0}=\emptyset$. Suppose we have defined $a_{\alpha}$ where $\alpha<\kappa$. Let $C^{\prime}_{\lambda\alpha}=\bigcap_{\beta<\alpha}C_{\sigma(\beta)}$; note that $C^{\prime}_{\lambda\alpha}\subseteq\kappa$ is club. Let $\eta\in C^{\prime}_{\lambda\alpha}$ be least such that $\eta\geq\max(\alpha,\mathrm{strsup}(a_{\alpha}))$. Then let $a^{\prime}_{\alpha+1}=a_{\alpha}\cup(\eta\oplus\pi(\alpha))$. Let $\beta$ be the least element of $B$ such that $\beta>\sup(a^{\prime}_{\alpha+1})$, and let $a_{\alpha+1}=a^{\prime}_{\alpha+1}\cup\{\beta\}$.
Now suppose we have defined $a_{\alpha}$ for all $\alpha<\zeta$, where $\zeta<\kappa$ is a limit. Let $a^{\prime}_{\zeta}=\bigcup_{\alpha<\zeta}a_{\alpha}$. Then set $a_{\zeta}=a^{\prime}_{\zeta}\cup\{\beta\}$, where $\beta$ is least such that $\beta\in B$ and $\beta\geq\sup(a^{\prime}_{\zeta})+\zeta$. (Note the “$+\zeta$” here introduces “extra space”; this is needed to see that we get the club $C_{\lambda}$ with the desired properties.) Now set $A_{\lambda}=\bigcup_{\alpha<\kappa}a_{\alpha}$. Also set $C_{\lambda}$ to be the set of all limit ordinals $\alpha$ such that $\alpha=\sup(a^{\prime}_{\alpha})$.
Clearly $A_{\lambda}\cap B$ is unbounded in $\kappa$, and properties (i), (ii) and (iv) hold for all $\alpha\leq\lambda$. For (iii), fix $\beta<\lambda$; we verify that $A_{\beta}\cap A_{\lambda}$ is bounded in $\kappa$. Well, $A_{\beta}\cap B$ is bounded by choice of $B$. All elements of $A_{\lambda}\backslash B$ are in $a^{\prime}_{\alpha+1}\backslash a_{\alpha}$ for some $\alpha<\kappa$. But for all sufficiently large $\alpha<\kappa$, we have $\beta\in\sigma``\alpha$, so $C^{\prime}_{\lambda\alpha}\subseteq C_{\beta}$, so letting $\eta$ be as in the definition of $a^{\prime}_{\alpha+1}$, we have $\eta\in C_{\beta}$, so
$A_{\beta}\cap(a^{\prime}_{\alpha+1}\backslash a_{\alpha})\subseteq A_{\beta}\cap[\eta,\eta+\eta)=\emptyset,$
which suffices.
It is straightforward to see the construction does not stop at a stage $\lambda<\kappa^{+}$, and =$\mathscr{F}=\{A_{\alpha}\}_{\alpha<\kappa^{+}}$ is a mad family of cardinality $\kappa^{+}$.
We now want to see that $\mathscr{F}$ is $\Pi_{1}(\{\kappa,\mathfrak{m}\})$. Well, given $A\subseteq\kappa$, we have $A\in\mathscr{F}$ iff the following two conditions hold:
  1. 1.
    Every bounded subset of $\kappa$ is directly encoded into $A$.
  2. 2.
    For all premice $P$ such that $\mathrm{OR}^{P}=\kappa$ and $A\cap\alpha\in P$ for all $\alpha<\kappa$ and $\mathfrak{m}=P|\omega_{1}^{P}$ and $P\models$“I am the condensation stack over $\mathfrak{m}$ (see [19])”, and all sound premice $Q$ such that $P\triangleleft Q$, $\rho_{\omega}^{Q}=\kappa$, $Q\models$$\kappa^{+}$ exists”, and $Q$ satisfies the condensation theorem with respect to elementary $\pi:\bar{Q}\to Q$ with $\mathrm{cr}(\pi)=\rho_{\omega}^{\bar{Q}}$ and $\pi(\mathrm{cr}(\pi))=\kappa$ (see [25]), if $A\in Q$ then $Q\models$“there is $\alpha<\kappa^{+}$ such that $A=A_{\alpha}$” (where we define ($A_{\alpha}\right>_{\alpha<\kappa^{+Q}})^{Q}$ over $Q$ from $\mathfrak{m},\kappa$ just as $A_{\alpha}\right>_{\alpha<\kappa^{+}}$ was defined over $M$ from $\mathfrak{m},\kappa$ above).
This equivalence is straightforward to verify. Since these conditions are $\Pi_{1}(\{\mathfrak{m},\kappa\})$, this completes the proof of part 1.
Part 2: As $M$ is tame, [16] shows that, working in $M$, there is $N\triangleleft M|\omega_{1}$ such that $\mathfrak{m}$ is the unique premouse $P$ with universe $\mathrm{HC}$, such that $N\triangleleft P$ and $P$ is above-$\mathrm{OR}^{N}$, $(\omega,\omega_{1})$-iterable. We will use $x=N$ in our definition of $\mathscr{F}$. It follows that for each $A\subseteq\kappa$, we have $A\in\mathscr{F}$ iff the following conditions hold:
  1. 1.
    Every bounded subset of $\kappa$ is directly encoded into $A$.
  2. 2.
    For every premouse $P$ with $\mathrm{OR}^{P}=\kappa$ and $\{A\cap\alpha\bigm{|}\alpha<\kappa\}\subseteq P$ and $N\triangleleft P$, and every above-$N$, $(\omega,\omega_{1}^{P})$-strategy for $P|\omega_{1}^{P}$ (using $\mathrm{HC}^{P}=\mathrm{HC}$ to determine the countable iteration trees involved), if $P\models$“I am the condensation stack…” (as in the $\Pi_{1}(\{\kappa,\mathfrak{m}\})$ definition of $\mathscr{F}$ in part 1), then for all $Q$ with $P\triangleleft Q$ and as in part 1, if $A\in Q|\kappa^{+Q}$ then $A=A_{\alpha}^{Q}$ for some $\alpha<\kappa^{+Q}$.
Part 3: We just give a sketch here. In case $M=M_{n}$, then $\mathfrak{m}$ is definable over $\mathrm{HC}=\mathrm{HC}^{M}$, which, combined with the foregoing arguments, suffices. If $M$ is the minimal proper class mouse with a Woodin limit of Woodins, then we use that $\mathfrak{m}$ is $(\omega,\omega_{1}+1)$-iterable in $M$.  [^10] This is a standard fact, but is non-trivial. One can compute the relevant Q-structure for determining the correct $\mathcal{T}$-cofinal branch through an $\omega$-maximal tree $\mathcal{T}$ on $\mathfrak{m}$ by a kind of background construction. Extenders $E$ overlapping $M(\mathcal{T})$ can be added by deriving them from core embeddings, when a model $N_{\alpha}$ of the construction projects $<\delta(\mathcal{T})$, setting $N_{\alpha+1}=(N_{\alpha}||\gamma,E)$ for the appropriate cardinal $\gamma$ of $N_{\alpha}$. One does this until reaching some $N_{\alpha}$ such that the $\delta(\mathcal{T})$-core of $N_{\alpha}$ is the desired Q-structure for $M(\mathcal{T})$ (this might also project $<\delta(\mathcal{T})$). This is enough to define $\mathfrak{m}$ over $\mathcal{H}_{\kappa}$ if $\kappa>\omega_{1}$. If $\kappa=\omega_{1}$ we need to do some more work, because with $\Pi_{1}(\{\kappa\})$ we can then only directly assert that the premouse $P$ under consideration (with universe $\mathrm{HC}$) is $(\omega,\omega_{1})$-iterable. We claim that given this, it follows that $P=\mathfrak{m}$. The proof of this uses some of the ideas from [16], but here there is less going on, so we explain how it works in the present context. Suppose $P\neq\mathfrak{m}$, and let $\Gamma$ be an $(\omega,\omega_{1})$-strategy for $P$, and $\Sigma$ an $(\omega,\omega_{1}+1)$-strategy for $\mathfrak{m}$. Compare the pair $(\mathfrak{m},P)$ using $(\Sigma,\Gamma)$, producing a pair $(\mathcal{T},\mathcal{U})$ of trees, each of length $\omega_{1}$. Let $b=\Sigma(\mathcal{T})$. Let $Q\trianglelefteq M^{\mathcal{T}}_{b}$ be the Q-structure for $\delta(\mathcal{T})=\omega_{1}$, so $Q$ determines $b$ in the usual manner. We claim that $Q$ similarly determines a $\mathcal{U}$-cofinal branch, which gives the usual contradiction. For let $\pi:W\to\mathcal{H}_{\omega_{2}}$ be $\Sigma_{10}$-elementary, where $W$ is countable and transitive, with $(\mathcal{T},\mathcal{U}),b\in\mathrm{rg}(\pi)$. Let $\pi(\bar{Q},\bar{b},\bar{\mathcal{T}},\bar{\mathcal{U}},\eta)=(Q,b,\mathcal{T},\mathcal{U},\omega_{1})$. Then $[0,\eta)_{\mathcal{T}}=\bar{b}=b\cap\eta$ is determined by $\bar{Q}$. Let $R=M(\mathcal{T})=M(\mathcal{U})$. Suppose first that $R\models$“There is a proper class of Woodins”. Then by the minimality of $M$, $Q=\mathcal{J}_{\alpha}(R)$ for some $\alpha\in\mathrm{OR}$, so $\bar{Q}=\mathcal{J}_{\bar{\alpha}}(R|\eta)$ for some $\bar{\alpha}$. It follows that $\bar{Q}\triangleleft R$ and $\bar{Q}$ is a (hence the) Q-structure for $\eta$ in $R$, but then $\bar{Q}$ did determine $[0,\eta)_{\mathcal{U}}$, as desired. Now suppose instead that $R\models$“There is not a proper class of Woodins”. Then $\delta<\eta$, where $\delta$ is the sup of Woodins in $R$. We have $\bar{Q}\trianglelefteq M^{\mathcal{T}}_{\eta}$, but since $\eta$ is not Woodin in $R$, it follows that $\mathrm{OR}^{\bar{Q}}<\min(\mathrm{lh}(E^{\mathcal{T}}_{\eta}),\mathrm{lh}(E^{\mathcal{U}}_{\eta}))$. But then $\bar{Q}\triangleleft R=M(\mathcal{U})$ and $\bar{Q}\triangleleft M^{\mathcal{U}}_{\eta}$, so $\bar{Q}$ does determine $[0,\eta)_{\mathcal{U}}$, as desired. ∎
8 Independent families
8.1 Definition.
For $\kappa$ an uncountable cardinal, an independent family at $\kappa$ is a set $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ such that for all finite sets $\mathscr{A},\mathscr{B}\subseteq\mathscr{F}$ with $\mathscr{A}\cap\mathscr{B}=\emptyset$, letting =$\mathscr{B}^{\prime}=\{\kappa\backslash B\bigm{|}B\in\mathscr{B}\}$, we have
$\mathrm{card}\Big{(}\big{(}\bigcap\mathscr{A}\big{)}\cap\big{(}\bigcap\mathscr{B}^{\prime}\big{)}\Big{)}=\kappa.$
A maximal independent family is an independent family $\mathscr{F}$ for which there is no independent family $\mathscr{F}^{\prime}$ with $\mathscr{F}\subsetneq\mathscr{F}^{\prime}$.
There is a result for maximal independent families analogous to that for mad families given by Theorem 7.2:
8.2 Theorem.
Let $M$ be a $(0,\omega_{1}+1)$-iterable premouse, and let $\kappa\in\mathrm{OR}^{M}$ be such that $M\models$$\kappa>\omega$ is a regular cardinal and $\kappa^{+}$ exists”. Let $\mathfrak{m}=M|\omega_{1}^{M}$. Then:
  1. 1.
    $M\models$“there is a maximal independent family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa,\mathfrak{m}\})$”.
  2. 2.
    If $M$ is tame then $M\models$“there is a maximal independent family $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa,x\})$ for some $x\in\mathbb{R}$”.
  3. 3.
    If $M=M_{n}$ for some $n<\omega$ (the minimal proper class mouse with $n$ Woodin cardinals) or $M$ is the minimal proper class mouse satisfying “there is a Woodin limit of Woodin cardinals” then $M\models$“there is a maximal independent familiy $\mathscr{F}\subseteq\mathcal{P}(\kappa)$ of cardinality $\kappa^{+}$ which is $\Pi_{1}(\{\kappa\})$”.
Proof.
We use a method similar to that for mad families. The key is again to arrange that for every $A\in\mathscr{F}$, every bounded subset of $\kappa$ is directly encoded into $A$. But the methods for instantiating this are somewhat different to those we used for mad families.
Work in $M$. Fix a canonical surjection $\pi:\kappa\to\mathcal{P}({<\kappa})$ with $\pi(\alpha)\subseteq\alpha$ for each $\alpha<\kappa$, as in the proof of Theorem 7.2. We construct a sequence $\mathscr{F}_{\alpha}\right>_{\alpha\leq\kappa^{+}}$ of independent families $\mathscr{F}_{\alpha}\subseteq\mathcal{P}(\kappa)$, recursively in $\alpha$, maintaining that each $\mathscr{F}_{\alpha}$ is an independent family of cardinality $\kappa$, and for every $A\in\mathscr{F}_{\alpha}$, every bounded subset of $\kappa$ is directly encoded into $A$, and such that $\mathscr{F}_{\alpha}\subseteq\mathscr{F}_{\beta}$ for all $\alpha<\beta<\kappa^{+}$. At limit stages $\lambda$ we just take a union, setting $\mathscr{F}_{\lambda}=\bigcup_{\alpha<\lambda}\mathscr{F}_{\alpha}$; note that this maintains the inductive hypotheses. At the end, defining $\mathscr{F}=\mathscr{F}_{\kappa^{+}}$, this will be a maximal independent family.
Now let us define $\mathscr{F}_{0}$. Let $I_{\alpha}\right>_{\alpha<\kappa}$ be the least enumeration of pairwise disjoint intervals $\subseteq\kappa$, such that each $I_{\alpha}$ has length $\alpha$. Say that $\alpha<\kappa$ is even if it has form $\eta+2n$ for some ordinal $\eta$ which is either 00 or a limit ordinal and some $n<\omega$, and odd otherwise. Let $\mathrm{OR}_{\mathrm{even}}$ be the class of even ordinals, and $\mathrm{OR}_{\mathrm{odd}}$ the odd ordinals. Let $B_{0}=\bigcup_{\alpha\in\kappa\cap\mathrm{OR}_{\mathrm{even}}}I_{\alpha}$ and $B_{1}=\bigcup_{\alpha\in\kappa\cap\mathrm{OR}_{\mathrm{odd}}}I_{\alpha}$. Let $B_{0}^{\prime}$ be the least subset of $B_{0}$ such that every bounded subset of $\kappa$ is directly encoded as a subset of $I_{\alpha}$ for some even $\alpha$. Let $\mathscr{G}$ be the $M$-least independent family at $\kappa$ of cardinality $\kappa$ with $\bigcup\mathscr{G}\subseteq B_{1}$. Now set
$\mathscr{F}_{0}=\{B_{0}^{\prime}\cup A\bigm{|}A\in\mathscr{G}\}.$
Note that $\mathscr{F}_{0}$ is an independent family of cardinality $\kappa$, and every $A\in\mathscr{F}_{0}$ directly encodes every bounded subset of $\kappa$.
Now suppose we have defined $\mathscr{F}_{\alpha}$ with the right properties; it remains to construct $\mathscr{F}_{\alpha+1}$. Let $A\subseteq\kappa$ be $M$-least such that $A\notin\mathscr{F}_{\alpha}$ and $\mathscr{F}_{\alpha}\cup\{A\}$ is independent, if such an $A$ exists. We will define $\mathscr{F}_{\alpha+1}$ to be (an independent family) such that if $A\notin\mathscr{F}_{\alpha+1}$ then $\mathscr{F}_{\alpha+1}\cup\{A\}$ is not independent, maintaining the inductive hypotheses. We will do this by adding exactly two new sets, $B_{0}$ and $B_{1}$ (we will set $\mathscr{F}_{\alpha+1}=\mathscr{F}_{\alpha}\cup\{B_{0},B_{1}\}$), with $A\subseteq B_{0}\cup B_{1}$, which certainly takes care of making $\mathscr{F}_{\alpha+1}\cup\{A\}$ non-independent, if $A\notin\{B_{0},B_{1}\}$. For the purposes of the definability of $\mathscr{F}$, we will also ensure that $A$ itself can easily be computed from each of $B_{0}$ and $B_{1}$.
Let $\sigma:\kappa\to(\mathscr{F}_{\alpha})^{<\omega}\times(\mathscr{F}_{\alpha})^{<\omega}$ be the $M$-least function for each $\beta<\kappa$, if $\sigma(\beta)=(X,Y)$ then $X\cap Y=\emptyset$, and such that $\sigma^{-1}(C)$ has cardinality $\kappa$ for each $(X,Y)\in(\mathscr{F}_{\alpha})^{<\omega}\times(\mathscr{F}_{\alpha})^{<\omega}$ such that $X\cap Y=\emptyset$. Let
$\tau(\beta)=\Big{(}\bigcap_{X\in\mathrm{rg}(\sigma(\beta)_{0})}X\Big{)}\cap\Big{(}\bigcap_{X\in\mathrm{rg}(\sigma(\beta)_{1}}(\kappa\backslash X)\Big{)}$
where $(x,y)_{0}=x$ and $(x,y)_{1}=1$. So $\tau(\beta)\cap A$ and $\tau(\beta)\backslash A$ are both unbounded in $\kappa$ for each $\beta$ (by choice of $A$). We construct $B_{0},B_{1}$ through $\kappa$-many stages $\beta<\kappa$ by recursion on $\beta$, determining some $\eta_{\beta}<\kappa$ with $\beta\leq\eta_{\beta}$, and $B_{0}\cap\eta_{\beta}$ and $B_{1}\cap\eta_{\beta}$, before stage $\beta$. Set $\eta_{0}=0$. Consider stage $\beta$, given $\eta_{\beta}$, $B_{0}\cap\eta_{\beta}$ and $B_{1}\cap\eta_{\beta}$. Let us now specify $B_{0}\cap[\eta_{\beta},\omega\eta_{\beta})$ and $B_{1}\cap[\eta_{\beta},\omega\eta_{\beta})$.
We will first arrange that $\pi(\beta)$ will be directly encoded into $B_{0}$ and into $B_{1}$. For $n<\omega$, we first put
$[(2n+1)\eta_{\beta},(2n+2)\eta_{\beta})\subseteq B_{1},$
$[(2n+2)\eta_{\beta},(2n+3)\eta_{\beta})\subseteq B_{0}.$
Note that this ensures that $[\eta_{\beta},\omega\eta_{\beta})\subseteq B_{0}\cup B_{1}$, so that we certainly won’t violate the requirement that $A\subseteq B_{0}\cup B_{1}$ within the interval $[\eta_{\beta},\omega\eta_{\beta})$. Now we directly encode $\pi(\beta)$ in $B_{0}$ at $\eta_{\beta}$ (and since $\pi(\beta)\subseteq\beta\leq\eta_{\beta}$, this requires at most the interval $[\eta_{\beta},2\eta_{\beta})$; we don’t put any further ordinals into $B_{0}\cap[\eta_{\beta},2\eta_{\beta})$ beyond those needed for this). And we also directly encode $\pi(\beta)$ in $B_{1}$ at $2\eta_{\beta}$ (within the interval $[2\eta_{\beta},3\eta_{\beta})$). We now want to ensure that $A\cap[\eta_{\beta},\omega\eta_{\beta})$ can be easily recovered from $B_{i}\cap[3\eta_{\beta},\omega\eta_{\beta})$, for $i\in\{0,1\}$. We do this by sliding the sets $A_{[(n+1)\eta_{\beta},(n+2)\eta_{\beta})}$ up appropriately, staggered between $B_{0}$ and $B_{1}$, into the remaining intervals we have available. For $\eta\leq\delta$, let $A_{[\eta,\delta)}=\{\xi\bigm{|}\eta+\xi\in A\cap\delta\}$. Then we directly encode $A_{[\eta_{\beta},2\eta_{\beta})}$ in $B_{0}$ at $3\eta_{\beta}$ (which determines $B_{0}\cap[3\eta_{\beta},4\eta_{\beta})$) and in $B_{1}$ at $4\eta_{\beta}$ (determining $B_{1}\cap[4\eta_{\beta},5\eta_{\beta})$). Likewise, we directly encode $A_{[(n+1)\eta_{\beta},(n+2)\eta_{\beta})}$ in $B_{0}$ at $(3+2n)\eta_{\beta}$ and in $B_{1}$ and $(4+2n)\eta_{\beta}$, for each $n<\omega$. This determines $B_{0}\cap[\eta_{\beta},\omega\eta_{\beta})$ and $B_{1}\cap[\eta_{\beta},\omega\eta_{\beta})$.
We next want to take measures to help insure the independence of $\mathscr{F}_{\alpha}\cup\{B_{0},B_{1}\}$. We find the least 3 ordinals $\xi_{0}<\xi_{1}<\xi_{2}$ such that $\xi_{i}\geq\omega\eta_{\beta}$ and $\xi_{0},\xi_{1},\xi_{2}\in\tau(\beta)$, and then the least ordinal $\xi_{3}$ such that $\xi_{2}<\xi_{3}$ and $\xi_{3}\in\tau(\beta)\backslash A$. We then put all ordinals in the interval $[\omega\eta_{\beta},\xi_{3}]$ into both $B_{0},B_{1}$, excepting that we declare $\xi_{1},\xi_{3}\notin B_{0}$ and $\xi_{2},\xi_{3}\notin B_{1}$. This determines $B_{0}\cap(\xi_{3}+1)$ and $B_{1}\cap(\xi_{3}+1)$. Let $\gamma$ be such that $\omega\eta_{\beta}+\gamma=\xi_{3}+1$. We then set
$[\omega\eta_{\beta}+(2n+1)\gamma,\omega\eta_{\beta}+(2n+2)\gamma)\subseteq B_{1},$
$[\omega\eta_{\beta}+(2n+2)\gamma,\omega\eta_{\beta}+(2n+3)\gamma)\subseteq B_{0}.$
We then, much like before, directly encode $A_{[\omega\eta_{\beta}+n\gamma,\omega\eta_{\beta}+(n+1)\gamma)}$ in $B_{0}$ at $\omega\eta_{\beta}+(2n+1)\gamma$, and in $B_{1}$ at $\omega\eta_{\beta}+(2n+2)\gamma$. Defining $\eta_{\beta+1}=\omega\eta_{\beta}+\omega\gamma$, this determines $B_{0}\cap\eta_{\beta+1}$ and $B_{1}\cap\eta_{\beta+1}$, completing this round. Note that $A\cap[\eta_{\beta},\eta_{\beta+1})\subseteq(B_{0}\cup B_{1})\cap[\eta_{\beta},\eta_{\beta+1})$, since in fact $\xi_{3}\notin A$ and
$[\eta_{\beta},\eta_{\beta+1})\cap(B_{0}\cup B_{1})=[\eta_{\beta},\eta_{\beta+1})\backslash\{\xi_{3}\}.$
At limits $\lambda<\kappa$ set $\eta_{\lambda}=\sup_{\alpha<\lambda}\eta_{\alpha}$.
This completes the construction of $B_{0},B_{1}$. Clearly each bounded subset of $\kappa$ is encoded directly into both $B_{0},B_{1}$, and $A\subseteq B_{0}\cup B_{1}$. And note that every finite Boolean combination of elements $\mathscr{F}_{\alpha}\cup\{B_{0},B_{1}\}$ is unbounded in $\kappa$ (recall that $\sigma^{-1}(X,Y)$ is unbounded in $\kappa$ for each $(X,Y)\in(\mathscr{F}_{\alpha})^{<\omega}\times(\mathscr{F}_{\alpha})^{<\omega}$ with $X\cap Y=\emptyset$). So we have maintained the inductive requirements.
Now we want to see that $\mathscr{F}$ is appropriately definable. This is much like in the proof of Theorem 7.2, using the following observation. Given $B\subseteq\kappa$, we have $B\in\mathscr{F}$ iff for all $Q\triangleleft M$ with $\rho_{\omega}^{Q}=\kappa$ and $Q\models$$\kappa^{+}$ exists” and $B\in Q$, we have $B\in\mathscr{F}^{Q}$, where $\mathscr{F}^{Q}$ is defined over $Q$ just as $\mathscr{F}$ was over $M$. For it’s easy to see that $\mathscr{F}^{Q}\subseteq\mathscr{F}$ (in fact $\mathscr{F}^{Q}_{\alpha}=\mathscr{F}_{\alpha}$ for all $\alpha<\kappa^{+Q}$). And conversely, if $B\in\mathscr{F}$ then fix $\alpha<\kappa^{+}$ such that $B\in\{B_{0},B_{1}\}$ where these are the sets produced at stage $\alpha$, and let $A$ be the corresponding set at that stage. Then because $A$ is easily recovered from $B$, we have $A\in Q|\kappa^{+Q}$, and it follows that $\alpha<\kappa^{+Q}$, and so $B\in\mathscr{F}_{\alpha+1}=\mathscr{F}_{\alpha+1}^{Q}$. Using this observation, the methods used in the proof of Theorem 7.2 allow us to produce an appropriate definition of $\mathscr{F}$. ∎
9 Filters above measurables
We make a couple of simple observations on definability of filters.
9.1 Theorem.
Assume ZFC, $\kappa$ is a limit of measurable cardinals and $\operatorname{cof}(\kappa)=\omega$. Then there is no $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ ultrafilter over $\kappa$ which contains no bounded subsets of $\kappa$.
Proof.
After arranging for fixing the defining elements, just consider a sequence of measurables $\kappa_{0}<\kappa_{1}<\ldots$ cofinal in $\kappa$, and the sets $X=\cup_{n<\omega}[\kappa_{2n},\kappa_{2n+1})$ and $Y=\bigcup_{n<\omega}[\kappa_{2n+1},\kappa_{2n+2})$. Either $X\in U$ or $Y\in U$. Say $X\in U$ (otherwise it is similar). Then we can easily iterate at the $\kappa_{n}$’s, sending $X$ to some $X^{\prime}$ with $X\cap X^{\prime}=\emptyset$ (and we have chosen the $\kappa_{n}$’s so that the relevant iteration will fix the defining parameters). This is a contradiction. ∎
Let $\mathrm{Club}_{\kappa}$ denote the club filter at $\kappa$. It was shown by Sy-David Friedman and Liuzhen Wu in [3, Proposition 2.1] that if $\kappa$ is weakly compact then $\mathrm{Club}_{\kappa}$ is not $\Pi_{1}(\mathcal{H}_{\kappa^{+}})$. And it was shown by Philipp Lücke, Ralf Schindler and Philipp Schlicht in [10, Theorem 1.9] that if $\kappa$ is a regular cardinal which is a stationary limit of $\omega_{1}$-iterable cardinals then $\mathrm{Club}_{\kappa}$ is not $\Pi_{1}(\{\kappa\})$. We prove a variant of these results here. Note that neither part of Theorem 9.2 is a direct consequence of the results just mentioned; in particular in part 2, we do not demand that $\kappa$ be a stationary limit of measurables (and we also allow arbitrary parameters in $V_{\kappa}\cup\mathrm{OR}$). If $\kappa$ is the least inaccessible limit of measurables, then the set $C$ of all limits of measurables $<\kappa$ is club in $\kappa$ and consists of singular cardinals. But every $\omega_{1}$-iterable cardinal is inaccessible (see [5, Propositions 2.8, 3.1]), so [10, Theorem 1.9] requires that $\kappa$ be a stationary limit of inaccessibles (and more).
9.2 Theorem.
Let $\kappa$ be an uncountable cardinal. Then:
  1. 1.
    Suppose $\mu<\kappa$ and $\mu$ is measurable, $\kappa$ is $\mu$-steady and $\operatorname{cof}(\kappa)>\mu$. Then $\mathrm{Club}_{\kappa}$ is not $\Pi_{1}(V_{\mu}\cup\{\kappa,\kappa^{+}\})$. In fact, let $U$ be any normal measure on $\mu$ and let $S_{j}=\{x\bigm{|}j(x)=x\}$ where $j=i^{V}_{U}:V\to M=\mathrm{Ult}(V,U)$ is the ultrapower map (so $V_{\mu}\cup\{\kappa,\kappa^{+}\}\subseteq S_{j}$ ). Then $\mathrm{Club}_{\kappa}$ is not $\Pi_{1}(S_{j})$.
  2. 2.
    Suppose $\kappa$ is an inaccessible limit of measurables. Then $\mathrm{Club}_{\kappa}$ is not $\Pi_{1}(V_{\kappa}\cup\mathrm{OR})$.
Proof.
Part 1: Suppose otherwise. Then the set of stationary subsets of $\kappa$ is $\Sigma_{1}(S_{j})$, since given $S\subseteq\kappa$, $S$ is stationary iff $\kappa\backslash S\notin\mathrm{Club}_{\kappa}$. Since this equivalence is preserved by $j$ and $j(\kappa)=\kappa$, we have that for each $S\in\mathcal{P}(\kappa)\cap\mathrm{Ult}(V,U)$, if $\mathrm{Ult}(V,U)\models$$S$ is stationary” then $S$ is stationary. Let $S=\{\alpha<\kappa\bigm{|}\operatorname{cof}^{\mathrm{Ult}(V,U)}(\alpha)=\mu^{+}\}$. Then $\mathrm{Ult}(V,U)\models$$S$ is stationary” (using the fact that $\operatorname{cof}(\kappa)>\mu$). But $S$ is not stationary, a contradiction. For let $C^{-}=j``\kappa$ and let $C$ be the closure of $C^{-}$. So $C$ is club in $\kappa$, so it suffices to see that $C\cap S=\emptyset$. But $S\cap\mathrm{rg}(j)=\emptyset$, and every $\alpha\in C\backslash C^{-}$ is such that $\operatorname{cof}(\alpha)=\mu$, and since $\mathrm{Ult}(V,U)$ is closed under $\mu$-sequences, therefore $\operatorname{cof}^{\mathrm{Ult}(V,U)}(\alpha)=\mu$ also, so $\alpha\notin S$. This completes the proof.
Part 2: This is an immediate corollary of part 1 and Lemma 2.11. ∎
10 Global regularity properties
In this section we describe a variant of [15, Theorem 2.19]. The adaptation involves standard forcing techniques; see for example [2].
10.1 Theorem.
Assume ZFC + $\kappa$ is $\lambda$-supercompact, where $\lambda>\kappa$ is inaccessible and $2^{\lambda}=\lambda^{+}$. Then in a forcing extension $V[G]$, $\kappa$ is $\lambda$-supercompact, $(2^{\kappa})^{V[G]}=\kappa^{+V[G]}=\lambda$ and for every $X\subseteq\mathcal{P}(\kappa)^{V[G]}$ such that $X$ is definable over $V[G]$ from elements of $V\cup\mathcal{P}(\kappa)^{V[G]}$ and $X$ has cardinality $>\kappa$, we have that $X$ has a perfect subset, and $X$ is not a wellorder.
Proof.
Let $\mathbb{P}$ be length $\kappa$ Easton support iteration where at inaccessible stages $\alpha$, we force with the Levy collapse $\mathrm{Coll}(\alpha,{<\lambda_{\alpha}})$, where $\lambda_{\alpha}$ is the least inaccessible $>\alpha$, and at other stages $\alpha$ the empty forcing. Let +=*$\mathbb{P}^{+}=\mathbb{P}*\dot{\mathrm{Coll}}(\kappa,{<\lambda})$. Let $G^{+}$ be $(V,\mathbb{P}^{+})$-generic and $G=G^{+}\!\upharpoonright\!\kappa$ (so $G$ is $(V,\mathbb{P})$-generic).
Let $j:V\to M$ be a $\lambda$-supercompactness embedding. Note that $\mathbb{P}^{+}\in M$ and $G^{+}$ is $(M,\mathbb{P}^{+})$-generic. We have $\lambda=\kappa^{+V[G^{+}]}=\kappa^{+M[G^{+}]}$. Now $V\models$${{}^{\lambda}}M\subseteq M$”, and considering $\mathbb{P}^{+}$-names in $V$ for $\lambda$-sequences in $V[G^{+}]$, a straightforward calculation gives that $V[G^{+}]\models$(${{}^{\lambda}}(M[G^{+}])\subseteq M[G^{+}]$”.
Let $H^{+}=j``G^{+}$. So $H^{+}\in M[G^{+}]$. Working in $M[G^{+}]$, let $\widetilde{q}$ be the canonical $j(\mathbb{P})$-name for the union of those conditions of form $(j(p)_{j(\kappa)})_{G^{+}*\dot{I}}$ for $p\in G^{+}$, where $\dot{I}$ is the canonical name for the generic filter for the tail forcing $j(\mathbb{P})\!\upharpoonright\!(\kappa+1,j(\kappa))$, and where $j(p)_{\alpha}$ is the $\alpha$th component of $j(p)$. Note that in $M[G^{+}]$, $\widetilde{q}$ is forced (by $j(\mathbb{P})\!\upharpoonright\!(\kappa+1,j(\kappa))$) to be a $\mathrm{Col}(j(\kappa),{<j(\lambda)})$-name, because the collection of conditions is forced to be pairwise compatible, even closed under witnesses for such compatibility, and because the collection has size $\lambda$, which is smaller than the closure of the forcing (which is $j(\kappa)$). So in $M[G^{+}]$, $\widetilde{q}$ yields a condition in $j(\mathbb{P})\!\upharpoonright\!(\kappa+1,j(\kappa))*\mathrm{Coll}(j(\kappa),{<j(\lambda)})$.
Now working in $V[G^{+}]$, we construct an $M[G^{+}]$-generic filter $I$ for this tail forcing with $\widetilde{q}\in I$. This is possible because the forcing in question has size $j(\lambda)$ in $M[G^{+}]$ (which is inaccessible in $M[G^{+}]$), and $\mathrm{card}^{V[G^{+}]}(j(\lambda))=\kappa^{+V[G^{+}]}=\lambda$, so since $V\models$$2^{\lambda}=\lambda^{+}$”, there are only $\kappa^{++V[G^{+}]}=\lambda^{+V[G^{+}]}=\lambda^{+V}$- many dense subsets to meet, and the forcing is closed under $\lambda$-sequences in both $V[G^{+}]$ and $M[G^{+}]$, and $M[G^{+}]$ is closed under $\lambda$-sequences in $V[G^{+}]$. (In $V[G^{+}]$, enumerate those dense sets in ordertype $\lambda^{+V[G^{+}]}$, and meet them in turn.)
So $M[G^{+},I]\subseteq V[G^{+}]$. But now $j:V\to M$ extends to $j^{+}:V[G^{+}]\to M[G^{+},I]$ with $j^{+}(G^{+})=G^{+}*I$, and note that $j^{+}$ witnesses that $\kappa$ is $\kappa^{+V[G^{+}]}$-supercompact in $V[G^{+}]$.
It remains to verify the regularity properties of sets $X$, as claimed in the theorem. So suppose that $X\subseteq\mathcal{P}(\kappa)$ is definable over $V[G^{+}]$ from some $A\subseteq\kappa$ and some $a\in V$, and $X$ has cardinality $\lambda=\kappa^{+V[G^{+}]}$ in $V[G^{+}]$.
The fact that $X$ has a perfect subset follows from [15, Theorem 2.19], since $\kappa$ is regular in $V[G]$, $\lambda$ inaccessible in $V[G]$, and the last factor of $\mathbb{P}^{+}$ is just $\mathrm{Col}(\kappa,{<\lambda})^{V[G]}$.
Let us see that $X$ is not a wellorder. Suppose otherwise. There is some $\beta<\lambda$ such that $A\in V[G,(G^{+})_{\kappa}\!\upharpoonright\!\beta]$. Since $X$ has cardinality $>\kappa$, there must be some $B$ in the field of $X$ which is not in $V[G,(G^{+})_{\kappa}\!\upharpoonright\!\beta]$. Let $\xi\in\mathrm{OR}$ be the rank of $B$ in the wellorder. Then by the homogeneity of the tail forcing adding $(G^{+})_{\kappa}\!\upharpoonright\![\beta,\lambda)$, one can define $B$ from the parameters $A,a,\xi$ over $V[G,(G^{+})_{\kappa}\!\upharpoonright\!\beta]$, by consulting the forcing relation. So $B\in V[G,(G^{+})_{\kappa}\!\upharpoonright\!\beta]$, a contradiction. ∎
11 Questions
I now list some questions I would like to know the answers to. (ZFC is still the background.)
  1. 1.
    Let $\mu$ is measurable, and $U$ a $\mu$-complete ultrafilter over $\mu$. Suppose $\mu<\kappa$ and $\kappa$ is a $\mu$-steady cardinal, and $\kappa$ is not a limit of measurables. Can there be an ultrafilter $\mathscr{F}$ over $\kappa$ which contains all unbounded subsets of $\kappa$, and which is $\Sigma_{1}(V_{\mu}\cup S_{\kappa})$-definable, where $S_{\kappa}=\{x\bigm{|}i^{\mathcal{T}_{U}}_{0\alpha}(x)=x\text{ for all }\alpha<\kappa\}$? What if $\mathrm{Club}_{\kappa}\subseteq\mathscr{F}$? What if $V=L[U]$? What about “$\Pi_{1}$” replacing “$\Sigma_{1}$”?
  2. 2.
    Assume $\kappa$ is a limit of measurables. Can there be a $\Sigma_{1}(\mathcal{H}_{\kappa}\cup\mathrm{OR})$ maximal independent family $\subseteq\mathcal{P}(\kappa)$?
  3. 3.
    Does Theorem 7.2 generalize to the case that $\kappa$ is singular? What about Theorem 8.2?
  4. 4.
    Assume $V=L[U]$ where $U$ is a normal measure on $\mu$. Let $\kappa>\mu^{+}$ be $\mu$-steady. Let $X$ be as in the proof of Theorem 4.3 part 4. Is $X$ $\Pi_{1}(\{\kappa\})$-definable? It does not seem clear to the author that $A\in X$ iff for every $\kappa$-good filter $D$ and every ordinal $\alpha>\kappa$, if $A\in L_{\alpha}[D]$ then there is a limit ordinal $\eta<\kappa$ such that $L_{\alpha}[D]\models$$A\in\bigcap_{\beta<\eta}M^{\mathcal{T}_{D}}_{\beta}$ but $A\notin M^{\mathcal{T}_{D}}_{\eta}$”. For it seems that we might have $A\in X$ and some $\kappa$-good filter $D$ and some $\alpha<\kappa^{+}$ such that $A\in L_{\alpha}[D]$ but $A\notin\mathrm{Ult}(L_{\alpha}[D],D)$; in this case, since $A\in X$, we must actually have $A\in\mathrm{Ult}(L[D],D)$; it is just that $\alpha$ is not large enough to see where $A$ is constructed in $\mathrm{Ult}(L[D],D)$. But the author doesn’t know whether this actually occurs, and whether one could compute any useful bound on where in $\mathrm{Ult}(L[D],D)$ a set $A\subseteq\kappa$ can be constructed, given where it is constructed in $L[D]$.